### v1

# ASCoT: An Adaptive Self-Correction Chain-of-Thought Method for Late-Stage Fragility in LLMs

Abstract
Chain-of-Thought (CoT) prompting has significantly advanced the reasoning capabilities of Large Language Models (LLMs), yet the reliability of these reasoning chains remains a critical challenge. A widely held “cascading failure” hypothesis suggests that errors are most detrimental when they occur early in the reasoning process. This paper challenges that assumption through systematic error-injection experiments, revealing a counter-intuitive phenomenon we term “Late-Stage Fragility”: errors introduced in the later stages of a CoT chain are significantly more likely to corrupt the final answer than identical errors made at the beginning. To address this specific vulnerability, we introduce the Adaptive Self-Correction Chain-of-Thought (ASCoT) method. ASCoT employs a modular pipeline in which an Adaptive Verification Manager (AVM) operates first, followed by the Multi-Perspective Self-Correction Engine (MSCE). The AVM leverages a Positional Impact Score function $I(k)$ that assigns different weights based on the position within the reasoning chains, addressing the Late-Stage Fragility issue by identifying and prioritizing high-risk, late-stage steps. Once these critical steps are identified, the MSCE applies robust, dual-path correction specifically to the failure parts. Extensive experiments on benchmarks such as GSM8K and MATH demonstrate that ASCoT achieves outstanding accuracy, outperforming strong baselines, including standard CoT. Our work underscores the importance of diagnosing specific failure modes in LLM reasoning and advocates for a shift from uniform verification strategies to adaptive, vulnerability-aware correction mechanisms.
Introduction
The advent of Chain-of-Thought (CoT) has endowed Large Language Models (LLMs) with remarkable multi-step reasoning capabilities (Nye et al. 2021; Kojima et al. 2022), enabling them to deconstruct complex problems into a sequence of manageable sub-problems (Yao et al. 2023; Shinn et al. 2023). Recent studies, such as the o1 from OpenAI  (Jaech et al. 2024) and DeepSeekR1  (Guo et al. 2025) have shown that extended CoT reasoning improves LLM performance. However, the success of CoT and its more verbose variants has introduced challenges: prohibitive computational costs from long reasoning chains (Chen et al. 2025) and, more critically, the inherent brittleness of the reasoning process itself (Zhang et al. 2024; Ke et al. 2025). A single flaw is often sufficient to invalidate an entire process of thought, raising significant concerns about the trustworthiness of these models (Chowdhury and Caragea 2025).
Figure 1: The top part (a) illustrates the standard CoT reasoning process, which often generates redundant outputs and lacks an effective self-correction mechanism when errors occur. The bottom part (b) depicts the ASCoT method, which incorporates a robust mechanism for self-correction, addressing potential reasoning flaws at each stage.
A widely held belief within the research community, often termed the “cascading failure” hypothesis, posits that the primary threat to CoT integrity lies in early-stage errors (Tyen et al. 2023). The logic is intuitive: an initial mistake in problem comprehension or a miscalculation will inevitably propagate through the logical dependency graph of the reasoning chain, contaminating all subsequent steps. Consequently, much of the research on improving CoT robustness has implicitly focused on ensuring the correctness of the initial steps through more robust prompting or post-hoc verification (Liu et al. 2024c; Yi, Wang, and Li 2025).
In this work, we directly investigate this foundational assumption, seeking to answer a critical question: Are all errors in a CoT created equal in their impact? To find out this, we conduct a systematic error-injection experiment, which introduces controlled symbolic and numeric errors at varying positions within correct reasoning chains. We then instruct the LLM to resume the reasoning process from the point at which the error occurred. Our findings are striking and counter-intuitive. We discover that errors introduced in later stages are more likely to lead to an incorrect final answer. We term this phenomenon “Late-Stage Fragility”. An early error often appears to trigger the model’s latent self-correction mechanisms, allowing it to recover and still reach the correct solution. However, as the model progresses steadily, its reasoning process gains a form of “semantic commitment”, reducing relevant information entropy and flexibility. At the later stages, the model is unable to evaluate the validity of the final calculations.
This empirical discovery suggests that a significant portion of efforts to secure CoT reasoning may be misdirected. As illustrated in Figure  1(a), a standard CoT process can be both inefficient due to redundancy and unreliable due to its inability to correct these late-stage errors. Building on a rethinking of how we evaluate and how we correct LLM reasoning, we introduce the ASCoT method. In contrast to standard methods, ASCoT is designed to address both challenges simultaneously. As shown in Figure  1(b), it first employs a semantic pruning stage to enhance efficiency, followed by a targeted, risk-aware correction mechanism that focuses computational resources precisely on the most vulnerable steps to ensure a correct final answer.
We evaluate ASCoT across various models, including the LLaMA-3.1-8B-Instruct and the Qwen2.5-Instruct series (3B, 7B, and 14B), using two widely recognized reasoning benchmarks: GSM8K and MATH.
To sum up, our contributions are summarized as follows:
  • As far as we know, this is the first to identify and quantify the “Late-Stage Fragility” phenomenon in CoT reasoning, demonstrating that errors in later steps are significantly more detrimental than those in earlier steps.
  • We introduce ASCoT, a novel method specifically designed to counteract this fragility by precisely identifying and robustly correcting the high-impact, late-stage errors that are most critical to reasoning accuracy.
  • Our extensive experiments validate the effectiveness of ASCoT. We demonstrate that ASCoT achieves outstanding accuracy on the GSM8k and MATH benchmarks, validating our approach of diagnosing the specific failure mode before engineering a targeted solution.
Preliminary
This section establishes the foundational concepts that motivate our work, offering a review of some effective methods for identifying token importance within CoT.
Efficient CoT in LLM Reasoning
A critical question for CoT efficiency is whether every token in different position contributes equally to the final answer (Xia et al. 2025). Several early approaches (Hou et al. 2022; Zhang et al. 2023; Chen et al. 2024a) measure importance based on the self-information or surprisal from a unidirectional language model:
$I_{1}(x_{i})=-\log P(x_{i}|x_{<i};\theta_{\mathcal{M}_{L}}),\quad$ (1)
where $x_{i}$ is a given token, $\theta_{\mathcal{M}_{L}}$ denotes the LLM used to compute the confidence and $P(\cdot\mid\cdot;\theta_{\mathcal{M}_{L}})$ is the conditional probability assigned by the causal LLMs. This metric suffers from known limitations, such as positional dependency, where tokens at the end of sentences often receive lower importance scores due to the nature of perplexity in LLMs.
To address this, more advanced methods like LLMLingua-2 (Pan et al. 2024) utilize a bidirectional language model (Jawahar, Sagot, and Seddah 2019; Devlin et al. 2019) (e.g., BERT-style) trained specifically for token classification. This approach measures importance by the predicted probability of a token being essential (Achiam et al. 2023), conditioned on the entire sequence:
$I_{2}(x_{i})=P(x_{i}|x_{<n};\theta_{\mathcal{M}_{L}}),\quad$ (2)
where $\theta_{\mathrm{\mathcal{M}_{L}}}$ represents the parameters of the trained bidirectional model. Such methods confirm that tokens within a CoT sequence are highly variable in importance. For example, tokens comprising mathematical equations or key entities (Ma et al. 2025) are identified as critical, whereas semantic connectors such as “so” and “since” contribute much less to the final outcome of the reasoning  (Hou et al. 2025). This highlights a significant potential for enhancing CoT efficiency by pruning these redundant tokens.
Error Propagation in Reasoning Process
Beyond efficiency, the primary challenge for CoT is reliability. A single error can invalidate an entire reasoning process (Guo et al. 2017; Tyen et al. 2023). The conventional understanding of this problem is rooted in the “cascading failure” hypothesis, which posits that errors are most damaging when they occur early in the reasoning chain (Jiang et al. 2025; Neumann and Singh 2025). According to this view, an early mistake propagates through the logical dependency graph, contaminating all subsequent steps and ensuring an incorrect outcome. In this study, our preliminary investigations reveal a counter-intuitive dynamic that challenges this assumption. Through systematic error-injection experiments, we observe that errors introduced in the later stages of a reasoning chain are significantly more detrimental to the final outcome than identical errors introduced at the beginning. We term this phenomenon “Late-Stage Fragility”.
Method
Figure 2: Illustration of ASCoT Pipeline. ASCoT generates the $CoT_{initial}$ from the target LLM, which is then compressed by the IRM module to a predefined ratio $\gamma$. The compressed CoT is verified by the AVM module, and if the confidence score exceeds a threshold $\tau$, the problematic steps are sent to the MSCE module for error correction. The corrected $CoT_{final}$ is then used for fine-tuning. Finally, ASCoT enables reasoning on new problems with the specified compression ratio $\gamma$.
In this section, we detail the architecture of our method. ASCoT generates the $CoT_{initial}$ based on the target LLM. The $CoT_{initial}$ is then processed through the IRM module, reducing the output to a fixed ratio $\gamma$. The remaining CoT is passed through the AVM module, which performs a rigorous verification process. If the model’s risk score for any CoT step exceeds a predefined threshold $\tau$, indicating a potential error, the corresponding CoT step is forwarded to the MSCE module. The MSCE module applies internal and external validation techniques to correct errors, resulting in the corrected $CoT_{final}$. Ultimately, the refined, non-redundant $CoT_{final}$ is used for fine-tuning. We illustrate the training and inference process in Figure  2, respectively.
Intelligent Routing Mechanism
The IRM serves as the initial processing stage, responsible for preliminary CoT pruning and compression. The core idea is inspired by methods TokenSkip (Xia et al. 2025), which recognize that not all tokens in a CoT are equally important for the final reasoning outcome. The mechanism operates by first generating a complete CoT path for a given problem. Then it employs a trained token importance model to assign a semantic importance score to each token in the chain. Based on a predefined compression ratio $\gamma$, tokens with scores below the corresponding threshold are pruned, resulting in a compressed CoT, namely the final answer.
Figure 3: The details of verification and self-correction. AVM and MSCE working together to correct a faulty step $t_{k}$. AVM computes $R(t_{k})$. MSCE then applies intrinsic and extrinsic correction to produce the final corrected $CoT_{final}$.
Adaptive Verification Manager
AVM dynamically identifies which steps are most likely to be incorrect and then computes a Risk Score ${R}(t_{k})$ for each step $t_{k}$ by integrating two key assessments: the step’s confidence assessment $Q(t_{k})$ and its positional impact $I(k)$.
Confidence Assessment Score
To provide a holistic measure of the single reasoning step’s quality, we first define a comprehensive quality score $Q(t_{k})$, which is a composite metric that evaluates a step $t_{k}$ across four distinct dimensions: Logical Validity $V(t_{k})$, Factual Support $G(t_{k})$, Semantic Clarity $C(t_{k})$ and Process Utility $U(t_{k})$.
Logical validity $V(t_{k})$ evaluates the internal logical consistency of step $t_{k}$. The score is a binary value that indicates if the step $t_{k}$ is derivable from its context $(P,S_{<k})$, determined by a verification query to the LLM $\mathcal{M}_{L}$:
$V(t_{k})=\mathbb{I}(o_{k}^{v}\in y),$ (3)
where $o_{k}^{v}$ is the model’s output of step $t_{k}$, $y$ is the set of valid responses and $\mathbb{I}(\cdot)$ is the indicator function.
Factual support $G(t_{k})$ measures the correctness of all $n$ verifiable steps:
$G(t_{k})=\prod_{i=1}^{n}T(a_{i}),$ (4)
where $T(a_{i})\rightarrow\{0,1\}$ is the verification function that evaluates whether the $i$-th action $a_{i}$ correctly invokes the appropriate external tool.
Semantic clarity $C(t_{k})$ quantifies the clarity of step $t_{k}$ based on the average conditional probability of its tokens $\{y_{1},...,y_{m}\}$, conditioned on the prior context:
$C(t_{k})=\frac{1}{m}\sum_{i=1}^{m}\log P(y_{i}|c_{<i}),$ (5)
where $y_{i}$ is the $i$-th token of step $t_{k}$, $m$ is the total number of tokens, $c_{<i}$ is the preceding context of $y_{i}$ and $P(\cdot\mid\cdot)$ is the conditional probability produced by the LLM.
Process utility $U(t_{k})$ measures the contribution of step $t_{k}$ to the final answer $\mathcal{A}^{*}$. It is defined as the information gain, measured by the increase in the probability of the model generating $\mathcal{A}^{*}$ when $t_{k}$ is incorporated into the context:
$U(t_{k})=\log P(\mathcal{A}^{*}|S_{\leq k})-\log P(\mathcal{A}^{*}|S_{<k}),$ (6)
where $\mathcal{A}^{*}$ is the ground‐truth answer, $S_{\leq k}$ is the full context including step $t_{k}$ and $S_{<k}$is the context excluding step $t_{k}$.
Based on these dimensions, we design the single-step comprehensive quality scoring function $Q(t_{k})$ to integrate the multi-dimensional evaluation into a quantitative score. The function’s core design principle is a veto mechanism, implemented through a multiplicative structure. Logical validity and factual support are foundational. If a step fails on either of these criteria, it should be considered invalid regardless of its other attributes. Other assessment criteria play an auxiliary role. Therefore, the function is defined as:
$Q(t_{k})=V(t_{k})\,G(t_{k})\,\Bigl{(}w_{c}\,\frac{C(t_{k})}{C_{\max}}+w_{u}\,U(t_{k})\Bigr{)},$ (7)
where $w_{c}$ and $w_{u}$ are hyperparameters that represent the weight of semantic clarity and the weight for low process utility. Figure  3 illustrates this evaluation process, where an initial CoT with a calculation error is analyzed.
Positional Impact Score
To model “Late-Stage Fragility”, we introduce the Position Impact Function $I(k)$, which quantifies the influence of position on the system’s stability as it situates the different-stage regime:
$I(k)=w_{a}\cdot e^{\alpha(k-1)},$ (8)
where $k$ is the step’s position, $w_{a}$ is a base impact coefficient, and $\alpha$ is a propagation rate. These parameters are empirically fitted through controlled error-injection experiments. The detailed discussion of error propagation via controlled injection can be found in the experiment section. This procedure involves systematically injecting single errors at different steps $k$ in a set of baseline correct reasoning chains and measuring the probability of final answer failure. The function $I(k)$ is then fitted to these empirical data points. Finally, AVM combines these factors into a single Risk Score:
${R}(t_{k})=I(k)\,\times\bigl{(}1-Q(t_{k})\bigr{)}.$ (9)
This score is then compared against a tunable system threshold $\tau$. If $R(t_{k})>\tau$, the step is flagged as high-risk and passed to the MSCE for correction.
Multi-Perspective Self-Correction Engine
Once a step $t_{k}$ is flagged by the AVM, the MSCE is activated to perform a robust and efficient correction. To counteract a model struggles to correct its own mistakes when using a single line of reasoning, the MSCE employs a dual-path correction strategy: Intrinsic Correction (Self-Reflection), the model is presented with the context up to step $s_{(k-1)}$ and its own flawed step $t_{k}$, and is explicitly prompted to review and amend it. Extrinsic Correction (Independent Generation), the model is given the context up to $s_{(k-1)}$ but without the flawed step $t_{k}$. It is then prompted to generate a new subsequent step from this “clean slate”.
Experiment
Experimental Setup
Models and Datasets
We primarily evaluate our method using the LLaMA-3.1-8B-Instruct (Ge et al. 2023) and Qwen2.5-Instruct series (3B, 7B, 14B) (Team 2024)models. Our evaluation leverages two notably standard math reasoning benchmarks: GSM8K (Cobbe et al. 2021), which consists of 8.5K grade school math problems requiring 2-8 reasoning steps, and MATH (Hendrycks et al. 2021), a more challenging dataset of 12.5K competition-level math problems. Given the significant computational cost associated with the MATH dataset, we follow recent research practices by evaluating our method on a carefully selected subset, MATH-500 (Lightman et al. 2023), which includes a representative sample of 500 problems.
Implementation Details
The ASCoT is implemented in PyTorch 1.13.0 and all experiments are conducted on a server with four NVIDIA GeForce RTX 3090 GPUs (24GB). For the IRM module, we use LLMLingua-2 (Pan et al. 2024) as the token importance metric to generate compressed CoT data for fine-tuning. To optimize the fine-tuning process, we adopt LoRA (Hu et al. 2022) for all fine-tuning procedures, setting the rank $r=8$ and scaling parameter $\alpha=16$ to ensure efficient training. Additional implementation details and configurations can be found in the Appendix.
Methods Ratio GSM8K MATH-500
Accuracy $\uparrow$ Tokens $\downarrow$ ActRatio Accuracy $\uparrow$ Tokens $\downarrow$ ActRatio
Original 1.0 86.2(0.0↓) 213.17 1.00 48.6(0.0↓) 502.60 1.00
Prompt 0.9 84.1(2.1↓) 226.37 1.06 48.6(0.0↓) 468.04 0.93
0.7 84.9(1.3↓) 209.39 0.98 48.4(0.4↓) 472.13 0.94
0.5 83.7(2.5↓) 188.82 0.89 47.8(0.4↓) 471.11 0.94
Truncation 0.9 70.2(26.0↓) 202.06 0.95 47.8(0.8↓) 440.33 0.88
0.7 25.9(60.3↓) 149.97 0.70 45.0(3.6↓) 386.89 0.77
0.5 7.0(79.2↓) 103.69 0.49 27.4(21.2↓) 283.70 0.56
ASCoT 1.0 86.9(0.7↑) 214.24 1.00 48.8(0.2↑) 478.80 0.95
0.9 86.1(0.1↓) 197.65 0.92 47.6(1.0↓) 427.15 0.84
0.8 85.1(1.1↓) 169.96 0.79 47.2(1.4↓) 376.72 0.74
0.7 83.7(2.5↓) 151.32 0.71 46.8(1.8↓) 334.26 0.67
0.6 81.6(4.6↓) 128.38 0.60 41.8(6.8↓) 309.69 0.61
0.5 79.5(6.7↓) 118.57 0.55 40.1(8.5↓) 288.94 0.57
Table 1: Experimental results of ASCoT on LLaMA-3.1-8B-Instruct. We report Accuracy, average CoT token count (Tokens) and actual compression ratio (ActRatio) for comparison. The compression ratio $\gamma$ is set to 0.51.00.5-1.0, respectively.
Evaluation Metrics
We evaluate the performance of our method using three primary metrics: Accuracy (%), which represents the percentage of problems for which the final answer is correct. Tokens, the average number of tokens generated in the reasoning chain, measuring computational cost and efficiency during the reasoning process. ActRation, the actual compression ratio of ASCoT, reflecting the real efficiency of the token compression process.
Main Results
We evaluate the end-to-end performance of ASCoT against two prevalent CoT length control techniques. The first is Prompt, appends an explicit instruction (e.g., “Please reduce your CoT output by 50%”) to the model’s input, thus guiding the LLM to generate a shorter reasoning trace. The second baseline, termed Truncation, imposes a fixed upper bound on the number of output tokens during inference.
The results of LLaMA-3.1-8B-Instruct are summarized in Table 1. As shown, prompt‐based reduction consistently fails to meet the target ratio: even at a nominal 0.5 budget it yields an actual compression ratio of 0.94 on MATH-500, with only marginal savings in CoT tokens. Truncation strictly enforces the specified ratio but at the cost of dramatic performance degradation: at a budget of 0.5, the accuracy on GSM8K plummets by 79.2 points and by 21.2 points on MATH-500, rendering it unsuitable for tasks requiring robust multistep reasoning. In contrast, ASCoT closely matches the specified budget (e.g., ActRatio = 0.55 at Ratio = 0.5) while maintaining robust reasoning performance. In a full budget (Ratio = 1.0), ASCoT marginally outperforms the baseline, achieving 86.9% on GSM8K (+0.7) and 48.8% on MATH-500 (+0.2). In at a 0.5 budget, it maintains respectable accuracies of 79.5% (–6.7) on GSM8K and 40.1% (–8.5) on MATH-500, respectively. These results demonstrate that ASCoT strikes a favorable balance between compression and accuracy, validating its effectiveness.
We also provide the results of Qwen2.5‐Instruct series models as Table 2. Across the three model scales (3B, 7B, and 14B), ASCoT consistently maintains original performance while achieving substantial CoT compression. For the 3B model, ASCoT at 0.9 compression incurs only a 0.5 point drop in accuracy, while reducing token usage by 16% (314.87 $\rightarrow$ 262.31). As the compression ratio decreases to 0.5, the accuracy degrades from 83.5% to 74.9%, but the token count falls sharply to 170.87(reducing 45.7%). Similar trends hold for the 7B and 14B variants: at Ratio = 0.9, both scales suffer less than a 0.5 point accuracy drop (for the 14B model, there is even a slight improvement in performance), yet token counts shrink by about 15%. Even under aggressive compression (ratio = 0.5), the 14B model retains more than 91.6% accuracy, representing only a 1.5 point reduction, while halving the original token budget.
Figure 4: Performance comparison of ASCoT with varying maximum length constraints. We adjust the maximum length budget when evaluating our method on MATH-500.
Scale Methods Ratio Accuracy Tokens ActRatio
3B Original 83.7(0.0↓) 314.87
ASCoT 1.0 83.5(0.2↓) 318.32 1.00
0.9 83.2(0.5↓) 262.31 0.82
0.8 82.0(1.7↓) 250.75 0.79
0.7 80.1(3.6↓) 233.15 0.73
0.6 77.6(6.1↓) 199.44 0.63
0.5 74.9(8.8↓) 170.87 0.54
7B Original 91.4(0.0↓) 297.83
ASCoT 1.0 91.8(0.4↑) 294.74 1.00
0.9 91.3(0.1↓) 251.71 0.85
0.8 90.4(1.0↓) 234.95 0.80
0.7 90.3(1.1↓) 216.73 0.74
0.6 88.2(3.2↓) 178.28 0.60
0.5 86.3(5.1↓) 150.98 0.51
14B Original 93.1(0.0↓) 313.11
ASCoT 1.0 93.3(0.2↑) 312.30 1.00
0.9 93.6(0.5↑) 269.76 0.86
0.8 93.5(0.4↑) 246.57 0.79
0.7 93.7(0.6↑) 221.17 0.71
0.6 92.8(0.3↓) 177.40 0.57
0.5 91.6(1.5↓) 158.21 0.51
Table 2: Experimental results on the Qwen2.5-Instruct series. We also provide Accuracy, Tokens and ActRatio.
Error Propagation via Controlled Injection
To quantitatively investigate the dynamics of error propagation in CoT reasoning, we conduct a systematic error-injection study to test the “cascading failure” hypothesis. Our experimental setup begins with a baseline of correct CoT solutions for problems from the GSM8K benchmark. For each correct reasoning chain, we first introduce two distinct classes of perturbations into the reasoning steps of correct CoT sequences. Then, we evaluate how these perturbations affect the overall reasoning process.
The injected errors are not manually added incorrect tokens, but simulated based on real symbolic and numerical mistakes observed during the model’s own generation. By replicating these naturally occurring failure patterns in a controlled manner, we enable systematic analysis of how the position of such errors affects the final reasoning outcome.
Numeric errors were introduced to simulate arithmetic miscalculations by slightly altering a single number in the reasoning process (e.g., adding or subtracting a fixed constant value). Symbolic errors were introduced to simulate logical inconsistencies by replacing a primary arithmetic operator with its inverse (e.g. + to -, $\times$ to $\div$). We preemptively set strict constraints to avoid division by zero.
Table 3 reports the impact of single‐step perturbations generated by Qwen2.5-Instruct 7B. When injecting a numerical error, the accuracy drops are relatively uniform across positions: such as a miscalculation at step 4 a drop of 61.70%. In contrast, symbolic errors induce more severe degradation, and this severity escalates with later injection points. Replacement of an operator in the final step (4/4) inflicts a reduction of 80.80%. These results confirm that CoT reasoning exhibits “late-stage fragility” particularly to logical perturbations. More details can be found in the appendix.
Analysis
Length Budget Analysis
To further investigate ASCoT’s efficiency, we analyze its performance under a fixed token budget. As illustrated in Figure  4, we assess our approach under two distinct length-budget settings. In the first, the maximum generation length is scaled by the compression ratio (0.5–1.0). While in the second, we retain the original LLM budget (e.g., 1024 tokens in MATH) regardless of compression. The results reveal that under the unchanged budget, our method consistently surpasses the baseline: accuracy rises from 42.6% to a peak of 49.4%. Notably, at compression ratios of 0.8, and 0.9, we observe absolute gains of 0.8 and 0.6 percentage points over the baseline in a smaller tokens, underscoring the potential of compressed CoT within the same budgetary constraints.
The Late-Stage Fragility of CoT
Our findings reveal a phenomenon that challenges the traditional cascading failure model. To illustrate this conceptually, Figure  5 contrasts the model’s ability to recover from an identical type of error when introduced at different stages of the reasoning process.
An early-stage error, while potentially disruptive, was often corrected or mitigated by the model’s subsequent reasoning steps. However, an identical error introduced in the final or penultimate step was significantly more likely to go uncorrected, directly leading to a wrong answer.
Our findings do not aim to refute the cascading failure hypothesis entirely, but rather to refine and contextualize it. Previous work suggests that early-stage errors in CoT reasoning tend to propagate forward, compromising subsequent steps and the final answer. However, our analysis reveals that the impact of reasoning errors is not uniform. Errors introduced near the conclusion of reasoning are notably much less likely to be corrected. This indicates that cascading failures do occur, but that late-stage fragility constitutes an additional failure mode not previously emphasized.
Figure 5: A illustration of “Late-Stage Fragility”. The model recovers from an early-stage calculation error, but fails when an identical type of error is introduced in the final step. The tokens that occur related errors are highlighted in orange.
Case Study: ASCoT in Action
To provide a clear and detailed qualitative understanding of our method, we present a case study, illustrated in Figure  3. The problem asks for the total number of apples Bob picked over nine days.
As shown in the Initial CoT panel, a standard model correctly structures the reasoning but commits a calculation error in Step 2, incorrectly stating that 125 * 8 = 1,005. This single subtle yet critical flaw propagates directly to the final step, resulting in an incorrect answer of 1,020.
The Evaluation panel demonstrates the intervention of AVM module. The AVM performs a multi-dimensional assessment of the flawed step. It correctly identifies the lack of Logical Validity (0/1) due to the arithmetic error. Simultaneously, it recognizes that the step’s Factual Groundedness and Semantic Clarity & Process Utility are sound. This evaluation allows the AVM to pinpoint the exact failure without discarding the entire step’s valid structure.
Because the step’s risk score exceeds the threshold, the MSCE is activated. As shown in the Final result panel, the MSCE provides the corrected calculation (125 * 8 = 1,000) using a dual-path correction mechanism. This targeted repair allows the model to obtain the correct final answer of 1,015.
All/Err Type Ori Acc Fin Acc Drop Ratio
2/2 numeric 98.85 38.93 59.92
3/3 96.19 35.72 60.47
4/4 95.75 34.05 61.70
2/2 symbolic 98.85 29.18 69.67
3/3 96.19 20.42 75.77
4/4 95.75 14.95 80.80
Table 3: Error-injection results on GSM8K, using Qwen2.5-Instruct 7B. “All/Err” denotes the total number of reasoning steps and the step which a fault was injected. “Type” indicates the category of the injected error (numeric or symbolic). “Ori Acc” is the model’s accuracy before error-injection. “Fin Acc” is the accuracy after injection. “Drop Ratio” quantifies the relative decrease in accuracy.
Related Work
LLM Reasoning & Efficient Thinking
The advent of CoT prompting marked a significant leap in the reasoning capabilities of LLMs (Marconato et al. 2023; Ding et al. 2024; Liu et al. 2024d). CoT decomposes complex problems into a sequence of manageable intermediate steps, dramatically improving performance on arithmetic, commonsense, and symbolic reasoning tasks (Team 2024).
These efficiency-focused methods can be broadly categorized: Prompt-based strategies (Xu et al. 2025) aim to guide the model toward more concise reasoning through specific instructions in the prompt itself, such as setting a token budget (Han et al. 2024). Inference-time dynamic reasoning (Ding et al. 2025) includes early-exit mechanisms that terminate the reasoning process once a certain confidence threshold is met, avoiding unnecessary computation. CoT Distillation (Wadhwa, Amir, and Wallace 2024; Wang et al. 2023) focuses on transferring the reasoning ability of a large “teacher” model to a smaller “student” model.
CoT Compression and Pruning
The verbosity of CoT (Wingate, Shoeybi, and Sorensen 2022), while beneficial for reasoning, leads to high inference latency and computational costs. This has spurred research into CoT compression and pruning (Chuang et al. 2024; Xuefei et al. 2023), which aims to eliminate redundant information without degrading reasoning accuracy. This issue is often termed the “overthinking phenomenon,” (Kumar et al. 2025; Chen et al. 2024b) where models generate unnecessarily long chains that can even introduce errors.
To address this bottleneck, researchers have explored various techniques. Token-level pruning (Fu et al. 2024; Ma, Fang, and Wang 2023), uses a semantic importance metric to identify and remove less critical tokens from the reasoning chain, training the model to learn shortcuts between essential concepts. Other approaches operate at a more structural level. Prune-on-Logic (Liu et al. 2025) transforms a CoT into a logic graph and prunes entire reasoning steps based on a loss importance score. Simpler methods, such as direct truncation of the reasoning chain (Wang et al. 2024), have also been shown to be surprisingly effective, suggesting that complete reasoning paths are not always necessary.
Self-correction Mechanism
The capacity for self-correction is a critical frontier (Pan et al. 2023) for enhancing the reliability of LLMs. The general paradigm (Kamoi et al. 2024; Yang et al. 2025) involves a multi-stage process: generating an initial response, producing feedback on that response, and then refining it.
Approaches diverge based on the source of the feedback. Intrinsic self-correction (Li et al. 2024; Liu et al. 2024a) relies on the model’s own capabilities to critique and improve its output without external tools (Cao, Wang, and Xiong 2025; Tan et al. 2025). However, a key challenge is that models often fail to identify their own errors, a phenomenon attributed to a lack of error-correction examples in training data (Liu et al. 2024b; Anantaprayoon, Kaneko, and Okazaki 2025). Extrinsic self-correction incorporates feedback from external sources. This involves using a verifier model to assess the correctness of each step (Kim et al. 2025), query external knowledge, or check the consistency across outputs generated from multiple perspectives.
Conclusion
In this work, we challenge the conventional “cascading failure” hypothesis in CoT reasoning. Through systematic error-injection experiments, we identify and quantify “Late-Stage Fragility”: a counter-intuitive phenomenon where errors introduced near the end of a reasoning chain are significantly more detrimental than those at the beginning. To counteract this specific vulnerability, we introduce the Adaptive self-correction Chain-of-Thought (ASCoT) method, which is a targeted system that uses an adaptive verification manager to precisely identify and a multi-perspective engine to robustly correct these high-impact, late-stage errors. Our extensive experiments on the GSM8K and MATH benchmarks demonstrate that ASCoT achieves outstanding accuracy, outperforming strong baselines including the computationally expensive self-consistency. These results highlight the effectiveness of ASCoT in addressing late-stage fragility, validating the importance of diagnosing specific failure modes to engineer more effective solutions. This research advocates for a shift from uniform verification to adaptive, position-aware strategies. We hope our investigations pave the way for future advancements in correction mechanisms and enhance the reliability, robustness, and efficiency of LLM reasoning.

### v3

@@
 ASCoT: An Adaptive Self-Correction Chain-of-Thought Method for Late-Stage Fragility in LLMs
-Abstract
+Abstract.
 Chain-of-Thought (CoT) prompting has significantly advanced the reasoning capabilities of Large Language Models (LLMs), yet the reliability of these reasoning chains remains a critical challenge.
@@
 ASCoT employs a modular pipeline in which an Adaptive Verification Manager (AVM) operates first, followed by the Multi-Perspective Self-Correction Engine (MSCE).
-The AVM leverages a Positional Impact Score function $I(k)$ that assigns different weights based on the position within the reasoning chains, addressing the Late-Stage Fragility issue by identifying and prioritizing high-risk, late-stage steps.
+The AVM leverages a Positional Impact Score function that assigns different weights based on the position within the reasoning chains, addressing the Late-Stage Fragility issue by identifying and prioritizing high-risk, late-stage steps.
 Once these critical steps are identified, the MSCE applies robust, dual-path correction specifically to the failure parts.
-Extensive experiments on benchmarks such as GSM8K and MATH demonstrate that ASCoT achieves outstanding accuracy, outperforming strong baselines, including standard CoT.
-Our work underscores the importance of diagnosing specific failure modes in LLM reasoning and advocates for a shift from uniform verification strategies to adaptive, vulnerability-aware correction mechanisms.
-Introduction
-The advent of Chain-of-Thought (CoT) has endowed Large Language Models (LLMs) with remarkable multi-step reasoning capabilities (Nye et al.
-2021; Kojima et al.
-2022), enabling them to deconstruct complex problems into a sequence of manageable sub-problems (Yao et al.
-2023; Shinn et al.
-2023).
-Recent studies, such as the o1 from OpenAI  (Jaech et al.
-2024) and DeepSeekR1  (Guo et al.
-2025) have shown that extended CoT reasoning improves LLM performance.
-However, the success of CoT and its more verbose variants has introduced challenges: prohibitive computational costs from long reasoning chains (Chen et al.
-2025) and, more critically, the inherent brittleness of the reasoning process itself (Zhang et al.
-2024; Ke et al.
-2025).
-A single flaw is often sufficient to invalidate an entire process of thought, raising significant concerns about the trustworthiness of these models (Chowdhury and Caragea 2025).
-Figure 1: The top part (a) illustrates the standard CoT reasoning process, which often generates redundant outputs and lacks an effective self-correction mechanism when errors occur.
+Extensive experiments on benchmarks demonstrate our superiority over previous methods.
+ASCoT reduces token usage by 21%/30% for the LLaMA-3.1-8B on GSM8K/MATH-500 benchmarks, with minimal accuracy drops of only 1.1% and 1.8%, achieving new sota performance.
+Key words and phrases:
+Large Language Models, Common-Sense Reasoning, Interpretability and Evaluation of NLP Models, Prompt Engineering
+1. Introduction
+Figure 1.
+The top part (a) illustrates the standard CoT reasoning process, which often generates redundant outputs and lacks an effective self-correction mechanism when errors occur.
 The bottom part (b) depicts the ASCoT method, which incorporates a robust mechanism for self-correction, addressing potential reasoning flaws at each stage.
-A widely held belief within the research community, often termed the “cascading failure” hypothesis, posits that the primary threat to CoT integrity lies in early-stage errors (Tyen et al.
-2023).
+The advent of Chain-of-Thought (CoT) has endowed Large Language Models (LLMs) with remarkable multi-step reasoning capabilities Kojima et al.
+(2022), enabling them to deconstruct complex problems into a sequence of manageable sub-problems Yao et al.
+(2023); Shinn et al.
+(2023).
+Recent studies, such as the o1 from OpenAI Jaech et al.
+(2024) and DeepSeekR1 Guo et al.
+(2025) have shown that extended CoT reasoning improves LLM performance.
+However, the success of CoT and its variants has introduced challenges: prohibitive computational costs from long reasoning chains Chen et al.
+(2025) and the inherent brittleness of the reasoning process itself Zhang et al.
+(2024); Ke et al.
+(2025).
+A single flaw is often sufficient to invalidate an entire process of thought, raising significant concerns about the trustworthiness of these models Chowdhury and Caragea (2025).
+A widely held belief within the research community, often termed the “cascading failure” hypothesis, posits that the primary threat to CoT integrity lies in early-stage errors Tyen et al.
+(2023).
 The logic is intuitive: an initial mistake in problem comprehension or a miscalculation will inevitably propagate through the logical dependency graph of the reasoning chain, contaminating all subsequent steps.
-Consequently, much of the research on improving CoT robustness has implicitly focused on ensuring the correctness of the initial steps through more robust prompting or post-hoc verification (Liu et al.
-2024c; Yi, Wang, and Li 2025).
 In this work, we directly investigate this foundational assumption, seeking to answer a critical question: Are all errors in a CoT created equal in their impact?
@@
 An early error often appears to trigger the model’s latent self-correction mechanisms, allowing it to recover and still reach the correct solution.
-However, as the model progresses steadily, its reasoning process gains a form of “semantic commitment”, reducing relevant information entropy and flexibility.
+As the model progresses steadily, its reasoning process gains a form of “semantic commitment”, reducing relevant information entropy.
 At the later stages, the model is unable to evaluate the validity of the final calculations.
 This empirical discovery suggests that a significant portion of efforts to secure CoT reasoning may be misdirected.
-As illustrated in Figure  1(a), a standard CoT process can be both inefficient due to redundancy and unreliable due to its inability to correct these late-stage errors.
+As illustrated in Figure 1(a), a standard CoT process can be both inefficient due to redundancy and unreliable due to its inability to correct these late-stage errors.
 Building on a rethinking of how we evaluate and how we correct LLM reasoning, we introduce the ASCoT method.
 In contrast to standard methods, ASCoT is designed to address both challenges simultaneously.
-As shown in Figure  1(b), it first employs a semantic pruning stage to enhance efficiency, followed by a targeted, risk-aware correction mechanism that focuses computational resources precisely on the most vulnerable steps to ensure a correct final answer.
-We evaluate ASCoT across various models, including the LLaMA-3.1-8B-Instruct and the Qwen2.5-Instruct series (3B, 7B, and 14B), using two widely recognized reasoning benchmarks: GSM8K and MATH.
+As shown in Figure 1(b), it first employs a semantic pruning stage to enhance efficiency, followed by a targeted, risk-aware correction mechanism that focuses computational resources precisely on the most vulnerable steps to ensure a correct final answer.
 To sum up, our contributions are summarized as follows:
@@
 We demonstrate that ASCoT achieves outstanding accuracy on the GSM8k and MATH benchmarks, validating our approach of diagnosing the specific failure mode before engineering a targeted solution.
-Preliminary
+2. Preliminary
 This section establishes the foundational concepts that motivate our work, offering a review of some effective methods for identifying token importance within CoT.
-Efficient CoT in LLM Reasoning
-A critical question for CoT efficiency is whether every token in different position contributes equally to the final answer (Xia et al.
-2025).
-Several early approaches (Hou et al.
-2022; Zhang et al.
-2023; Chen et al.
-2024a) measure importance based on the self-information or surprisal from a unidirectional language model:
+Efficient CoT in LLM Reasoning A critical question for CoT efficiency is whether every token in different position contributes equally to the final answer Xia et al.
+(2025).
+Several early approaches Hou et al.
+(2022); Zhang et al.
+(2023); Chen et al.
+(2024b) measure importance based on the self-information or surprisal from a unidirectional language model:
 $I_{1}(x_{i})=-\log P(x_{i}|x_{<i};\theta_{\mathcal{M}_{L}}),\quad$ (1)
@@
 This metric suffers from known limitations, such as positional dependency, where tokens at the end of sentences often receive lower importance scores due to the nature of perplexity in LLMs.
-To address this, more advanced methods like LLMLingua-2 (Pan et al.
-2024) utilize a bidirectional language model (Jawahar, Sagot, and Seddah 2019; Devlin et al.
-2019) (e.g., BERT-style) trained specifically for token classification.
-This approach measures importance by the predicted probability of a token being essential (Achiam et al.
-2023), conditioned on the entire sequence:
-$I_{2}(x_{i})=P(x_{i}|x_{<n};\theta_{\mathcal{M}_{L}}),\quad$ (2)
-where $\theta_{\mathrm{\mathcal{M}_{L}}}$ represents the parameters of the trained bidirectional model.
-Such methods confirm that tokens within a CoT sequence are highly variable in importance.
-For example, tokens comprising mathematical equations or key entities (Ma et al.
-2025) are identified as critical, whereas semantic connectors such as “so” and “since” contribute much less to the final outcome of the reasoning  (Hou et al.
-2025).
-This highlights a significant potential for enhancing CoT efficiency by pruning these redundant tokens.
-Error Propagation in Reasoning Process
-Beyond efficiency, the primary challenge for CoT is reliability.
-A single error can invalidate an entire reasoning process (Guo et al.
-2017; Tyen et al.
-2023).
-The conventional understanding of this problem is rooted in the “cascading failure” hypothesis, which posits that errors are most damaging when they occur early in the reasoning chain (Jiang et al.
-2025; Neumann and Singh 2025).
-According to this view, an early mistake propagates through the logical dependency graph, contaminating all subsequent steps and ensuring an incorrect outcome.
-In this study, our preliminary investigations reveal a counter-intuitive dynamic that challenges this assumption.
-Through systematic error-injection experiments, we observe that errors introduced in the later stages of a reasoning chain are significantly more detrimental to the final outcome than identical errors introduced at the beginning.
-We term this phenomenon “Late-Stage Fragility”.
-Method
-Figure 2: Illustration of ASCoT Pipeline.
-ASCoT generates the $CoT_{initial}$ from the target LLM, which is then compressed by the IRM module to a predefined ratio $\gamma$.
-The compressed CoT is verified by the AVM module, and if the confidence score exceeds a threshold $\tau$, the problematic steps are sent to the MSCE module for error correction.
+To address this, more advanced methods like LLMLingua-2 Pan et al.
+(2024) utilize a bidirectional language model Jawahar et al.
+(2019); Devlin et al.
+(2019) (e.g., BERT-style) to reframe the task as token classification.
+In this paradigm, the importance of a token $x_{i}$ is defined as the probability that a trained classifier assigns it an ”essential” label ($y_{i}=1$), conditioned on the entire sequence $x$:
+$I_{2}(x_{i})=P(y_{i}=1|x;\theta_{\mathcal{M}{cls}}),$ (2)
+where $\theta_{\mathcal{M}{cls}}$ represents the parameters of the trained bidirectional classifier.
+Such methods confirm that tokens within a CoT sequence are variable in importance.
+For example, tokens comprising mathematical equations or key entities Ma et al.
+(2025) are identified as critical, whereas semantic connectors such as so” and since” contribute less to the final outcome of the reasoning Hou et al.
+(2025).
+This highlights that CoT efficiency can be improved by pruning these redundant tokens.
+Error Propagation in Reasoning Process Beyond efficiency, the primary challenge for CoT is reliability.
+A single error can invalidate an entire reasoning process Guo et al.
+(2017); Tyen et al.
+(2023).
+The conventional understanding of this problem is rooted in the “cascading failure” hypothesis, which posits that errors are most damaging when they occur early in the reasoning chain Jiang et al.
+(2025).
+According to this view, an early mistake propagates through the logical dependency graph, contaminating all subsequent steps.
+In this study, through systematic error-injection experiments, we observe that errors introduced in the later stages of a reasoning chain are significantly more detrimental to the final outcome than identical errors introduced at the beginning.
+Figure 2.
+Illustration of ASCoT Pipeline.
+ASCoT generates the $CoT_{initial}$ from the target LLM, which is then compressed by IRM module to a predefined ratio $\gamma$.
+The compressed CoT is verified by AVM module, and if the confidence score exceeds a threshold $\tau$, the problematic steps are sent to MSCE module for error correction.
 The corrected $CoT_{final}$ is then used for fine-tuning.
 Finally, ASCoT enables reasoning on new problems with the specified compression ratio $\gamma$.
+Algorithm 1 ASCoT Mechanism
+Input: Initial chain-of-thought (CoT) $CoT_{initial}$
+Output: Corrected $CoT_{final}$
+[…]
+[…]
+3:for each step $t_{k}$ in $CoT_{pruned}$ do
+4:  $is\_flagged\leftarrow\mathrm{AVM}(context,t_{k})$
+5:  if is_flaggedis\_flagged then
+6:   $t_{k}^{\mathrm{corr}}\leftarrow\mathrm{MSCE}(context,t_{k})$
+7:   append $t_{k}^{\mathrm{corr}}$ to $CoT_{final}$
+8:   $context\leftarrow context\,\|\,t_{k}^{\mathrm{corr}}$
+9:  else
+10:   append $t_{k}$ to $CoT_{final}$
+11:   $context\leftarrow context\,\|\,t_{k}$
+12:  end if
+13:end for
+14:return $CoT_{final}$
+3. Method
 In this section, we detail the architecture of our method.
@@
 Ultimately, the refined, non-redundant $CoT_{final}$ is used for fine-tuning.
-We illustrate the training and inference process in Figure  2, respectively.
-Intelligent Routing Mechanism
+We illustrate the training and inference process in Figure 2, respectively.
+We also provide Pseudocode 1 for ASCoT.
+3.1. Intelligent Routing Mechanism
 The IRM serves as the initial processing stage, responsible for preliminary CoT pruning and compression.
-The core idea is inspired by methods TokenSkip (Xia et al.
-2025), which recognize that not all tokens in a CoT are equally important for the final reasoning outcome.
+The core idea is inspired by methods TokenSkip Xia et al.
+(2025), which recognize that not all tokens in a CoT are equally important for the final reasoning outcome.
 The mechanism operates by first generating a complete CoT path for a given problem.
@@
 Based on a predefined compression ratio $\gamma$, tokens with scores below the corresponding threshold are pruned, resulting in a compressed CoT, namely the final answer.
-Figure 3: The details of verification and self-correction.
-AVM and MSCE working together to correct a faulty step $t_{k}$.
-AVM computes $R(t_{k})$.
-MSCE then applies intrinsic and extrinsic correction to produce the final corrected $CoT_{final}$.
-Adaptive Verification Manager
+3.2. Adaptive Verification Manager
 AVM dynamically identifies which steps are most likely to be incorrect and then computes a Risk Score ${R}(t_{k})$ for each step $t_{k}$ by integrating two key assessments: the step’s confidence assessment $Q(t_{k})$ and its positional impact $I(k)$.
-Confidence Assessment Score
-To provide a holistic measure of the single reasoning step’s quality, we first define a comprehensive quality score $Q(t_{k})$, which is a composite metric that evaluates a step $t_{k}$ across four distinct dimensions: Logical Validity $V(t_{k})$, Factual Support $G(t_{k})$, Semantic Clarity $C(t_{k})$ and Process Utility $U(t_{k})$.
+Confidence Assessment Score To provide a holistic measure of the single reasoning step’s quality, we first define a comprehensive quality score $Q(t_{k})$, which is a composite metric that evaluates step $t_{k}$ across four dimensions: Logical Validity $V(t_{k})$, Factual Support $G(t_{k})$, Semantic Clarity $C(t_{k})$ and Process Utility $U(t_{k})$.
 Logical validity $V(t_{k})$ evaluates the internal logical consistency of step $t_{k}$.
@@
 where $o_{k}^{v}$ is the model’s output of step $t_{k}$, $y$ is the set of valid responses and $\mathbb{I}(\cdot)$ is the indicator function.
-Factual support $G(t_{k})$ measures the correctness of all $n$ verifiable steps:
+Factual support $G(t_{k})$ measures the correctness of all $n$ verifiable assertions within the step.
+For the mathematical reasoning benchmarks used in this work, verifiable assertions are arithmetic calculations.
+The function is defined as:
 $G(t_{k})=\prod_{i=1}^{n}T(a_{i}),$ (4)
-where $T(a_{i})\rightarrow\{0,1\}$ is the verification function that evaluates whether the $i$-th action $a_{i}$ correctly invokes the appropriate external tool.
-Semantic clarity $C(t_{k})$ quantifies the clarity of step $t_{k}$ based on the average conditional probability of its tokens $\{y_{1},...,y_{m}\}$, conditioned on the prior context:
-$C(t_{k})=\frac{1}{m}\sum_{i=1}^{m}\log P(y_{i}|c_{<i}),$ (5)
-where $y_{i}$ is the $i$-th token of step $t_{k}$, $m$ is the total number of tokens, $c_{<i}$ is the preceding context of $y_{i}$ and $P(\cdot\mid\cdot)$ is the conditional probability produced by the LLM.
-Process utility $U(t_{k})$ measures the contribution of step $t_{k}$ to the final answer $\mathcal{A}^{*}$.
-It is defined as the information gain, measured by the increase in the probability of the model generating $\mathcal{A}^{*}$ when $t_{k}$ is incorporated into the context:
+where $a_{i}$ is the $i$-th arithmetic operation in step $t_{k}$, and $T(a_{i})\rightarrow\{0,1\}$ is a verification function that executes the operation using a computational engine (e.g., a Python interpreter) to confirm its accuracy.
+Semantic clarity $C(t_{k})$ quantifies the coherence of step $t_{k}$.
+To ensure the score is well-calibrated and bounded within a $[0,1]$ range, we normalize the average conditional log-probability of its tokens $\{y_{1},...,y_{m}\}$ using a sigmoid function:
+$C(t_{k})=\sigma\left(\frac{1}{m}\sum_{i=1}^{m}\log P(y_{i}|c_{<i})\right),$ (5)
+where $\sigma(\cdot)$ is the sigmoid function, $y_{i}$ is the $i$-th token of step $t_{k}$, and $c_{<i}$ is the preceding context.
+Process utility $U(t_{k})$ measures the contribution of step $t_{k}$ to the final answer $\mathcal{A}^{*}$. It is defined as the information gain:
 $U(t_{k})=\log P(\mathcal{A}^{*}|S_{\leq k})-\log P(\mathcal{A}^{*}|S_{<k}),$ (6)
 where $\mathcal{A}^{*}$ is the ground‐truth answer, $S_{\leq k}$ is the full context including step $t_{k}$ and $S_{<k}$is the context excluding step $t_{k}$.
-Based on these dimensions, we design the single-step comprehensive quality scoring function $Q(t_{k})$ to integrate the multi-dimensional evaluation into a quantitative score.
-The function’s core design principle is a veto mechanism, implemented through a multiplicative structure.
-Logical validity and factual support are foundational.
-If a step fails on either of these criteria, it should be considered invalid regardless of its other attributes.
-Other assessment criteria play an auxiliary role.
-Therefore, the function is defined as:
-$Q(t_{k})=V(t_{k})\,G(t_{k})\,\Bigl{(}w_{c}\,\frac{C(t_{k})}{C_{\max}}+w_{u}\,U(t_{k})\Bigr{)},$ (7)
-where $w_{c}$ and $w_{u}$ are hyperparameters that represent the weight of semantic clarity and the weight for low process utility.
-Figure  3 illustrates this evaluation process, where an initial CoT with a calculation error is analyzed.
-Positional Impact Score
-To model “Late-Stage Fragility”, we introduce the Position Impact Function $I(k)$, which quantifies the influence of position on the system’s stability as it situates the different-stage regime:
-$I(k)=w_{a}\cdot e^{\alpha(k-1)},$ (8)
-where $k$ is the step’s position, $w_{a}$ is a base impact coefficient, and $\alpha$ is a propagation rate.
-These parameters are empirically fitted through controlled error-injection experiments.
-The detailed discussion of error propagation via controlled injection can be found in the experiment section.
-This procedure involves systematically injecting single errors at different steps $k$ in a set of baseline correct reasoning chains and measuring the probability of final answer failure.
+$U(t_{k})$ relies on the ground-truth answer and is used during the training phase.
+At inference time, this term is omitted.
+Based on these dimensions, we design the single-step comprehensive quality scoring function $Q(t_{k})$ with a veto mechanism for foundational criteria:
+$Q(t_{k})=V(t_{k})\cdot G(t_{k})\cdot\left(w_{c}\cdot C(t_{k})+w_{u}\cdot U(t_{k})\right),$ (7)
+where $w_{c}$ and $w_{u}$ are hyperparameters that balance the contribution of clarity and utility.
+Figure 3 illustrates this evaluation process, where an initial CoT with a calculation error is analyzed.
+Figure 3.
+The details of verification and self-correction.
+AVM and MSCE working together to correct a faulty step $t_{k}$.
+AVM computes $R(t_{k})$.
+MSCE then applies intrinsic and extrinsic correction to produce the final corrected $CoT_{final}$.
+Positional Impact Score To quantitatively model the “Late-Stage Fragility” phenomenon observed in our error-injection experiments (Section 4.3), we introduce the Position Impact Function $I(k)$.
+This function is an empirical model fitted to our experimental data, designed to capture the escalating impact of errors at later reasoning stages.
+We normalize this position function as:
+$I(k)=w_{a}\cdot e^{\alpha(k/K)},$ (8)
+where $k$ is the step’s index, $K$ is the total number of steps, $w_{a}$ is a base impact coefficient, and $\alpha$ is a propagation rate.
+These parameters are determined by fitting the function to our error-injection data.
+This procedure involves injecting single errors at different steps $k$ in a set of baseline correct reasoning chains and measuring the probability of final answer failure.
 The function $I(k)$ is then fitted to these empirical data points.
+Details of the analysis are available in Appendix.
 Finally, AVM combines these factors into a single Risk Score:
@@
 This score is then compared against a tunable system threshold $\tau$. If $R(t_{k})>\tau$, the step is flagged as high-risk and passed to the MSCE for correction.
-Multi-Perspective Self-Correction Engine
-Once a step $t_{k}$ is flagged by the AVM, the MSCE is activated to perform a robust and efficient correction.
-To counteract a model struggles to correct its own mistakes when using a single line of reasoning, the MSCE employs a dual-path correction strategy: Intrinsic Correction (Self-Reflection), the model is presented with the context up to step $s_{(k-1)}$ and its own flawed step $t_{k}$, and is explicitly prompted to review and amend it.
-Extrinsic Correction (Independent Generation), the model is given the context up to $s_{(k-1)}$ but without the flawed step $t_{k}$.
-It is then prompted to generate a new subsequent step from this “clean slate”.
-Experiment
-Experimental Setup
-Models and Datasets
-We primarily evaluate our method using the LLaMA-3.1-8B-Instruct (Ge et al.
-2023) and Qwen2.5-Instruct series (3B, 7B, 14B) (Team 2024)models.
-Our evaluation leverages two notably standard math reasoning benchmarks: GSM8K (Cobbe et al.
-2021), which consists of 8.5K grade school math problems requiring 2-8 reasoning steps, and MATH (Hendrycks et al.
-2021), a more challenging dataset of 12.5K competition-level math problems.
-Given the significant computational cost associated with the MATH dataset, we follow recent research practices by evaluating our method on a carefully selected subset, MATH-500 (Lightman et al.
-2023), which includes a representative sample of 500 problems.
-Implementation Details
-The ASCoT is implemented in PyTorch 1.13.0 and all experiments are conducted on a server with four NVIDIA GeForce RTX 3090 GPUs (24GB).
-For the IRM module, we use LLMLingua-2 (Pan et al.
-2024) as the token importance metric to generate compressed CoT data for fine-tuning.
-To optimize the fine-tuning process, we adopt LoRA (Hu et al.
-2022) for all fine-tuning procedures, setting the rank $r=8$ and scaling parameter $\alpha=16$ to ensure efficient training.
-Additional implementation details and configurations can be found in the Appendix.
+Table 1.
+Experimental results of ASCoT on LLaMA-3.1-8B-Instruct.
+We report Accuracy, average CoT token count (Tokens) and actual compression ratio (ActRatio) for comparison.
+The compression ratio $\gamma$ is set to 0.51.00.5-1.0, respectively.
 Methods Ratio GSM8K MATH-500
@@
 0.5 79.5(6.7↓) 118.57 0.55 40.1(8.5↓) 288.94 0.57
-Table 1: Experimental results of ASCoT on LLaMA-3.1-8B-Instruct.
-We report Accuracy, average CoT token count (Tokens) and actual compression ratio (ActRatio) for comparison.
-The compression ratio $\gamma$ is set to 0.51.00.5-1.0, respectively.
-Evaluation Metrics
-We evaluate the performance of our method using three primary metrics: Accuracy (%), which represents the percentage of problems for which the final answer is correct.
+3.3. Multi-Perspective Self-Correction Engine
+Once a step $t_{k}$ is flagged by the AVM, the MSCE is activated to perform a robust and efficient correction.
+To counteract a model struggles to correct its own mistakes when using a single line of reasoning, the MSCE employs a dual-path correction strategy: Intrinsic Correction (Self-Reflection), the model is presented with the context up to step $s_{(k-1)}$ and its own flawed step $t_{k}$, and is explicitly prompted to review and amend it.
+Extrinsic Correction (Independent Generation), the model is given the context up to $s_{(k-1)}$ but without the flawed step $t_{k}$.
+Upon generating two candidate corrections, the MSCE employs a selection mechanism to choose the superior one.
+We re-apply the Confidence Assessment Score function $Q(\cdot)$ (from Eq.
+7, with $U=0$) to both candidates.
+The candidate with the higher quality score is then integrated into the reasoning chain.
+4. Experiment
+4.1. Experimental Setup
+Models and Datasets We primarily evaluate our method using the LLaMA-3.1-8B-Instruct Ge et al.
+(2023) and Qwen2.5-Instruct series (3B, 7B, 14B) Team (2024)models.
+Our evaluation leverages two notably standard math reasoning benchmarks: GSM8K Cobbe et al.
+(2021), which consists of 8.5K grade school math problems requiring 2-8 reasoning steps, and MATH Hendrycks et al.
+(2021), a more challenging dataset of 12.5K competition-level math problems.
+Given the significant computational cost associated with the MATH dataset, we follow recent research practices by evaluating our method on a carefully selected subset, MATH-500 Lightman et al.
+(2023), which includes a representative sample of 500 problems.
+Implementation Details The ASCoT is implemented in PyTorch 1.13.0 and all experiments are conducted on a server with four NVIDIA GeForce RTX 3090 GPUs (24GB).
+For the IRM module, we use LLMLingua-2 Pan et al.
+(2024) as the token importance metric to generate compressed CoT data for fine-tuning.
+To optimize the fine-tuning process, we adopt LoRA Hu et al.
+(2022) for all fine-tuning procedures, setting the rank $r=8$ and scaling parameter $\alpha=16$ to ensure efficient training.
+Additional implementation details and configurations can be found in the Appendix.
+Evaluation Metrics We evaluate the performance of ASCoT using three primary metrics: Accuracy (%), which represents the percentage of problems for which the final answer is correct.
 Tokens, the average number of tokens generated in the reasoning chain, measuring computational cost and efficiency during the reasoning process.
-ActRation, the actual compression ratio of ASCoT, reflecting the real efficiency of the token compression process.
-Main Results
+ActRatio, the actual compression ratio of ASCoT, reflecting the real efficiency of the token compression process.
+4.2. Main Results
 We evaluate the end-to-end performance of ASCoT against two prevalent CoT length control techniques.
 The first is Prompt, appends an explicit instruction (e.g., “Please reduce your CoT output by 50%”) to the model’s input, thus guiding the LLM to generate a shorter reasoning trace.
-The second baseline, termed Truncation, imposes a fixed upper bound on the number of output tokens during inference.
+The second termed Truncation, imposes a fixed upper bound on the number of output tokens during inference.
 The results of LLaMA-3.1-8B-Instruct are summarized in Table 1.
@@
 Truncation strictly enforces the specified ratio but at the cost of dramatic performance degradation: at a budget of 0.5, the accuracy on GSM8K plummets by 79.2 points and by 21.2 points on MATH-500, rendering it unsuitable for tasks requiring robust multistep reasoning.
-In contrast, ASCoT closely matches the specified budget (e.g., ActRatio = 0.55 at Ratio = 0.5) while maintaining robust reasoning performance.
+In contrast, ASCoT closely matches the specified budget (e.g., ActRatio = 0.79 at Ratio = 0.8 on GSM8K and ActRatio = 0.61 at Ratio = 0.6 on MATH-500) while maintaining robust reasoning performance.
 In a full budget (Ratio = 1.0), ASCoT marginally outperforms the baseline, achieving 86.9% on GSM8K (+0.7) and 48.8% on MATH-500 (+0.2).
@@
 These results demonstrate that ASCoT strikes a favorable balance between compression and accuracy, validating its effectiveness.
+Figure 4.
+Performance of ASCoT on the Qwen2.5-Instruct Series.
+The results for the 3B, 7B, and 14B models are shown under various compression ratios and compared against the original, uncompressed baseline.
+Figure 5. Performance comparison of ASCoT with varying maximum length constraints. We adjust the maximum length budget when evaluating our method on MATH-500.
 We also provide the results of Qwen2.5‐Instruct series models as Table 2.
@@
 Even under aggressive compression (ratio = 0.5), the 14B model retains more than 91.6% accuracy, representing only a 1.5 point reduction, while halving the original token budget.
-Figure 4: Performance comparison of ASCoT with varying maximum length constraints. We adjust the maximum length budget when evaluating our method on MATH-500.
+The detailed experimental results can be found in the appendix.
+Table 2. Experimental results of the Qwen2.5-Instruct series evaluated on GSM8K. We also provide Accuracy, Tokens and ActRatio.
 Scale Methods Ratio Accuracy Tokens ActRatio
 3B Original 83.7(0.0↓) 314.87
-ASCoT 1.0 83.5(0.2↓) 318.32 1.00
-0.9 83.2(0.5↓) 262.31 0.82
+ASCoT 1.0 83.5(0.2↓) 318.32 1.01
+0.9 83.2(0.5↓) 262.31 0.83
 0.8 82.0(1.7↓) 250.75 0.79
-0.7 80.1(3.6↓) 233.15 0.73
+0.7 80.1(3.6↓) 233.15 0.74
 0.6 77.6(6.1↓) 199.44 0.63
@@
 7B Original 91.4(0.0↓) 297.83
-ASCoT 1.0 91.8(0.4↑) 294.74 1.00
+ASCoT 1.0 91.8(0.4↑) 294.74 0.99
 0.9 91.3(0.1↓) 251.71 0.85
-0.8 90.4(1.0↓) 234.95 0.80
-0.7 90.3(1.1↓) 216.73 0.74
+0.8 90.4(1.0↓) 234.95 0.79
+0.7 90.3(1.1↓) 216.73 0.73
 0.6 88.2(3.2↓) 178.28 0.60
@@
 14B Original 93.1(0.0↓) 313.11
-ASCoT 1.0 93.3(0.2↑) 312.30 1.00
+ASCoT 1.0 93.3(0.2↑) 312.30 0.99
 0.9 93.6(0.5↑) 269.76 0.86
@@
 0.5 91.6(1.5↓) 158.21 0.51
-Table 2: Experimental results on the Qwen2.5-Instruct series. We also provide Accuracy, Tokens and ActRatio.
-Error Propagation via Controlled Injection
-To quantitatively investigate the dynamics of error propagation in CoT reasoning, we conduct a systematic error-injection study to test the “cascading failure” hypothesis.
+Figure 4 demonstrates that ASCoT consistently achieves substantial CoT compression across all tested Qwen2.5-Instruct series models (3B, 7B, and 14B).
+For the 3B model, a moderate compression that reduces token usage by nearly 17% (from 314.87 to 262.31 tokens) incurs only a negligible 0.5-point drop in accuracy.
+This trend is further amplified in the larger models.
+With the 7B variant, a 15% token reduction results in a minimal 0.1-point accuracy drop, and a nearly 50% reduction in tokens still maintains accuracy above 86%.
+Notably, the 14B model showcases ASCoT’s exceptional ability to enhance efficiency.
+Even when halving the original token budget (a 50% reduction to 158.21 tokens), the accuracy drop is a mere 1.5 percentage points.
+These results validate ASCoT’s effectiveness in striking a favorable balance between computational efficiency and reasoning fidelity, demonstrating its robustness and scalability across different model sizes.
+4.3. Error Propagation via Controlled Injection
+To quantitatively investigate the dynamics of error propagation in CoT reasoning, we conducted a comprehensive and systematic error-injection study to rigorously test the “cascading failure” hypothesis.
 Our experimental setup begins with a baseline of correct CoT solutions for problems from the GSM8K benchmark.
-For each correct reasoning chain, we first introduce two distinct classes of perturbations into the reasoning steps of correct CoT sequences.
-Then, we evaluate how these perturbations affect the overall reasoning process.
+For each correct reasoning chain, we first introduce two distinct classes of perturbations into the reasoning steps of the correct CoT sequences.
+Then, we assess and evaluate how these perturbations affect the overall reasoning process.
 The injected errors are not manually added incorrect tokens, but simulated based on real symbolic and numerical mistakes observed during the model’s own generation.
-By replicating these naturally occurring failure patterns in a controlled manner, we enable systematic analysis of how the position of such errors affects the final reasoning outcome.
+By replicating these naturally occurring failure patterns in a controlled manner, we enable systematic analysis of how the position and type of such errors affects the final reasoning outcome.
 Numeric errors were introduced to simulate arithmetic miscalculations by slightly altering a single number in the reasoning process (e.g., adding or subtracting a fixed constant value).
@@
 Table 4 reports the impact of single‐step perturbations generated by Qwen2.5-Instruct 7B.
-When injecting a numerical error, the accuracy drops are relatively uniform across positions: such as a miscalculation at step 4 a drop of 61.70%.
-In contrast, symbolic errors induce more severe degradation, and this severity escalates with later injection points.
-Replacement of an operator in the final step (4/4) inflicts a reduction of 80.80%.
-These results confirm that CoT reasoning exhibits “late-stage fragility” particularly to logical perturbations.
-More details can be found in the appendix.
-Analysis
-Length Budget Analysis
-To further investigate ASCoT’s efficiency, we analyze its performance under a fixed token budget.
-As illustrated in Figure  4, we assess our approach under two distinct length-budget settings.
-In the first, the maximum generation length is scaled by the compression ratio (0.5–1.0).
-While in the second, we retain the original LLM budget (e.g., 1024 tokens in MATH) regardless of compression.
+For numeric errors, the accuracy drops gradually across positions, with a miscalculation at step 2 causing a decrease of 14.64%, at step 3 a 29.56% drop, and at step 4 a significant reduction of 51.69%.
+In contrast, symbolic errors induce more severe performance degradation, especially when injected later in the reasoning process.
+A symbolic error at step 2 results in a large drop of 43.97%,.
+However, injecting a symbolic error at step 3 (3/3) results in an accuracy reduction of 56.78%.
+These results emphasize the susceptibility of CoT reasoning to errors, particularly with symbolic mistakes, which cause a more pronounced decline in performance as the reasoning progresses.
+Table 3 reports initial accuracy before error injection.
+All the above data are from dataset GSM8K.
+Table 3. Accuracy for Different Inference Steps in the GSM8K Dataset, using Qwen2.5-Instruct 7B.
+Steps Correct Total Accuracy
+0 615 660 93.18
+1 1 1 100.00
+2 777 786 98.85
+3 2044 2125 96.19
+4 1936 2022 95.75
+5 1110 1180 94.07
+6 440 484 90.91
+7 117 137 85.40
+8 39 58 67.24
+9 4 7 57.14
+10 11 11 100.00
+11 2 2 100.00
+4.4. Analysis
+Length Budget Analysis To further investigate ASCoT’s efficiency, we analyze its performance under a fixed token budget.
+As illustrated in Figure 5, we assess our approach under two distinct length-budget settings.
+In the first, the maximum generation length is scaled by the compression ratio (ranging from 0.5 to 1.0).
+In the second, we retain the original LLM budget (e.g., 512 tokens in GSM8K and 1024 tokens in MATH) regardless of compression.
 The results reveal that under the unchanged budget, our method consistently surpasses the baseline: accuracy rises from 42.6% to a peak of 49.4%.
-Notably, at compression ratios of 0.8, and 0.9, we observe absolute gains of 0.8 and 0.6 percentage points over the baseline in a smaller tokens, underscoring the potential of compressed CoT within the same budgetary constraints.
-The Late-Stage Fragility of CoT
-Our findings reveal a phenomenon that challenges the traditional cascading failure model.
-To illustrate this conceptually, Figure  5 contrasts the model’s ability to recover from an identical type of error when introduced at different stages of the reasoning process.
+Notably, at compression ratios of 0.8 and 0.9, we observe absolute accuracy gains of 0.8 and 0.6 percentage points over the baseline in a smaller tokens, underscoring the potential of compressed CoT reasoning within the same budgetary constraints.
+The Late-Stage Fragility of CoT Our findings reveal a phenomenon that challenges the traditional cascading failure model.
+To illustrate this conceptually, Figure 6 contrasts the model’s ability to recover from an identical type of error when introduced at different stages of the reasoning process.
 An early-stage error, while potentially disruptive, was often corrected or mitigated by the model’s subsequent reasoning steps.
@@
 This indicates that cascading failures do occur, but that late-stage fragility constitutes an additional failure mode not previously emphasized.
-Figure 5: A illustration of “Late-Stage Fragility”.
+Figure 6.
+A illustration of “Late-Stage Fragility”.
 The model recovers from an early-stage calculation error, but fails when an identical type of error is introduced in the final step.
 The tokens that occur related errors are highlighted in orange.
-Case Study: ASCoT in Action
-To provide a clear and detailed qualitative understanding of our method, we present a case study, illustrated in Figure  3. The problem asks for the total number of apples Bob picked over nine days.
+Case Study To provide a clear and detailed qualitative understanding of our method, we present a case study, illustrated in Figure 3.
+The problem asks for the total number of apples Bob picked over nine days.
 As shown in the Initial CoT panel, a standard model correctly structures the reasoning but commits a calculation error in Step 2, incorrectly stating that 125 * 8 = 1,005.
@@
 This targeted repair allows the model to obtain the correct final answer of 1,015.
-All/Err Type Ori Acc Fin Acc Drop Ratio
-2/2 numeric 98.85 38.93 59.92
-3/3 96.19 35.72 60.47
-4/4 95.75 34.05 61.70
-2/2 symbolic 98.85 29.18 69.67
-3/3 96.19 20.42 75.77
-4/4 95.75 14.95 80.80
-Table 3: Error-injection results on GSM8K, using Qwen2.5-Instruct 7B.
+Figure 3 illustrates this evaluation process, where an initial CoT with a calculation error is analyzed.
+5. Related Work
+5.1. LLM Reasoning & Efficient Thinking
+The advent of CoT prompting marked a significant leap in the reasoning capabilities of LLMs Marconato et al.
+(2023); Ding et al.
+(2024); Liu et al.
+(2024a).
+CoT decomposes complex problems into a sequence of manageable intermediate steps, dramatically improving performance on arithmetic, commonsense, and symbolic reasoning tasks Team (2024).
+These efficiency-focused methods can be broadly categorized: Prompt-based strategies Xu et al.
+(2025) aim to guide the model toward more concise reasoning through specific instructions in the prompt itself, such as setting a token budget Han et al.
+(2024).
+Inference-time dynamic reasoning Ding et al.
+(2025) includes early-exit mechanisms that terminate the reasoning process once a certain confidence threshold is met, avoiding unnecessary computation.
+CoT Distillation Wadhwa et al.
+(2024); Wang et al.
+(2023) focuses on transferring the reasoning ability of a large “teacher” model to a smaller “student” model.
+Table 4.
+Error-injection results on GSM8K, using Qwen2.5-Instruct 7B.
 “All/Err” denotes the total number of reasoning steps and the step which a fault was injected.
-“Type” indicates the category of the injected error (numeric or symbolic).
+“Type” indicates the category of the injected error.
 “Ori Acc” is the model’s accuracy before error-injection.
 “Fin Acc” is the accuracy after injection.
-“Drop Ratio” quantifies the relative decrease in accuracy.
-Related Work
-LLM Reasoning & Efficient Thinking
-The advent of CoT prompting marked a significant leap in the reasoning capabilities of LLMs (Marconato et al.
-2023; Ding et al.
-2024; Liu et al.
-2024d).
-CoT decomposes complex problems into a sequence of manageable intermediate steps, dramatically improving performance on arithmetic, commonsense, and symbolic reasoning tasks (Team 2024).
-These efficiency-focused methods can be broadly categorized: Prompt-based strategies (Xu et al.
-2025) aim to guide the model toward more concise reasoning through specific instructions in the prompt itself, such as setting a token budget (Han et al.
-2024).
-Inference-time dynamic reasoning (Ding et al.
-2025) includes early-exit mechanisms that terminate the reasoning process once a certain confidence threshold is met, avoiding unnecessary computation.
-CoT Distillation (Wadhwa, Amir, and Wallace 2024; Wang et al.
-2023) focuses on transferring the reasoning ability of a large “teacher” model to a smaller “student” model.
-CoT Compression and Pruning
-The verbosity of CoT (Wingate, Shoeybi, and Sorensen 2022), while beneficial for reasoning, leads to high inference latency and computational costs.
-This has spurred research into CoT compression and pruning (Chuang et al.
-2024; Xuefei et al.
-2023), which aims to eliminate redundant information without degrading reasoning accuracy.
-This issue is often termed the “overthinking phenomenon,” (Kumar et al.
-2025; Chen et al.
-2024b) where models generate unnecessarily long chains that can even introduce errors.
+“Accu Drop(abs.
+points)” quantifies the relative decrease in accuracy.
+All/Err Type Ori Acc Fin Acc Accu Drop
+4/2 numeric 95.75 81.11 14.64
+4/3 95.75 66.19 29.56
+4/4 95.75 44.06 51.69
+2/2 symbolic 98.85 54.88 43.97
+3/2 96.19 78.28 17.91
+3/3 96.19 39.41 56.78
+5.2. CoT Compression and Pruning
+The verbosity of CoT Wingate et al.
+(2022), while beneficial for reasoning, leads to high inference latency and computational costs.
+This has spurred research into CoT compression and pruning Chuang et al.
+(2024); Xuefei et al.
+(2023), which aims to eliminate redundant information without degrading reasoning accuracy.
+This issue is often termed the “overthinking phenomenon,” Kumar et al.
+(2025); Chen et al.
+(2024a) where models generate unnecessarily long chains that can even introduce errors.
 To address this bottleneck, researchers have explored various techniques.
-Token-level pruning (Fu et al.
-2024; Ma, Fang, and Wang 2023), uses a semantic importance metric to identify and remove less critical tokens from the reasoning chain, training the model to learn shortcuts between essential concepts.
+Token-level pruning Fu et al.
+(2024); Ma et al.
+(2023), uses a semantic importance metric to identify and remove less critical tokens from the reasoning chain, training the model to learn shortcuts between essential concepts.
 Other approaches operate at a more structural level.
-Prune-on-Logic (Liu et al.
-2025) transforms a CoT into a logic graph and prunes entire reasoning steps based on a loss importance score.
-Simpler methods, such as direct truncation of the reasoning chain (Wang et al.
-2024), have also been shown to be surprisingly effective, suggesting that complete reasoning paths are not always necessary.
-Self-correction Mechanism
-The capacity for self-correction is a critical frontier (Pan et al.
-2023) for enhancing the reliability of LLMs.
-The general paradigm (Kamoi et al.
-2024; Yang et al.
-2025) involves a multi-stage process: generating an initial response, producing feedback on that response, and then refining it.
+Prune-on-Logic Liu et al.
+(2025) transforms a CoT into a logic graph and prunes entire reasoning steps based on a loss importance score.
+Simpler methods, such as direct truncation of the reasoning chain Wang et al.
+(2024), have also been shown to be surprisingly effective, suggesting that complete reasoning paths are not always necessary.
+5.3. Self-correction Mechanism
+The capacity for self-correction is a critical frontier Pan et al.
+(2023) for enhancing the reliability of LLMs.
+The general paradigm Kamoi et al.
+(2024); Yang et al.
+(2025) involves a multi-stage process: generating an initial response, producing feedback on that response, and then refining it.
 Approaches diverge based on the source of the feedback.
-Intrinsic self-correction (Li et al.
-2024; Liu et al.
-2024a) relies on the model’s own capabilities to critique and improve its output without external tools (Cao, Wang, and Xiong 2025; Tan et al.
-2025).
-However, a key challenge is that models often fail to identify their own errors, a phenomenon attributed to a lack of error-correction examples in training data (Liu et al.
-2024b; Anantaprayoon, Kaneko, and Okazaki 2025).
+Intrinsic self-correction Li et al.
+(2024); Liu et al.
+(2024c) relies on the model’s own capabilities to critique and improve its output without external tools Cao et al.
+(2025); Tan et al.
+(2025).
+However, a key challenge is that models often fail to identify their own errors, a phenomenon attributed to a lack of error-correction examples in training data Liu et al.
+(2024b); Anantaprayoon et al.
+(2025).
 Extrinsic self-correction incorporates feedback from external sources.
-This involves using a verifier model to assess the correctness of each step (Kim et al.
-2025), query external knowledge, or check the consistency across outputs generated from multiple perspectives.
-Conclusion
+This involves using a verifier model to assess the correctness of each step Kim et al.
+(2025), query external knowledge, or check the consistency across outputs generated from multiple perspectives.
+6. Limitations
+While ASCoT demonstrates promising performance, we acknowledge that there still exist several limitations that define avenues for future research.
+The efficacy of the framework is inherently dependent on the self-evaluation capabilities of its base Large Language Model (LLM); if a model cannot fundamentally recognize its own fallacies, the utility of the AVM and MSCE modules is consequently constrained.
+Furthermore, our current evaluation is focused on numeric and symbolic errors within the specific domain of mathematical reasoning.
+Therefore, the generalization of ASCoT to other error modalities—such as factual, commonsense, or complex logical errors—and its transferability to diverse long-chain reasoning tasks like code generation, legal analysis, or scientific summarization, remain important open questions requiring further study.
+7. Conclusion
 In this work, we challenge the conventional “cascading failure” hypothesis in CoT reasoning.
@@
 We hope our investigations pave the way for future advancements in correction mechanisms and enhance the reliability, robustness, and efficiency of LLM reasoning.
+Appendix A Appendix
+This document provides supplementary material, including pseudocode for the ASCoT mechanism, experimental setup, error-injection experiments, and results analysis.
+It also covers model configurations, hyperparameters used in the study.
+Additionally, the document discusses the model’s late-stage fragility and outlines current limitations and future directions.
+A.1. Experimental Setup
+To clearly elucidate the algorithmic pipeline of the ASCoT methodology, we provide its pseudocode.
+The algorithm commences upon receiving a problem query, which is first processed by the target Large Language Model (LLM) to generate an initial CoT solution, denoted as $CoT_{initial}$.
+This initial output is subsequently passed to the IRM module for token pruning based on a predefined compression ratio $\gamma$, thereby enhancing efficiency.
+The compressed CoT then undergoes a step-wise risk assessment by the AVM module.
+For each step $t_{k}$, the AVM computes a Risk Score $R(t_{k})$, which integrates a multi-dimensional quality assessment $Q(t_{k})$ with a Positional Impact score $I(k)$.
+If this score exceeds a predefined threshold $\tau$, the step is flagged as high-risk and is routed to the MSCE module for remediation.
+The MSCE employs a dual-path strategy, encompassing intrinsic reflection and extrinsic independent generation to correct the identified error, yielding the final, refined $CoT_{final}$.
+Model Configuration and Hyperparameters All experiments were conducted on an Ubuntu 20.04 LTS server equipped with four NVIDIA GeForce RTX 3090 GPUs (24GB each), using PyTorch 1.13.0.
+We fine‑tuned each base LLM via Low‑Rank Adaptation (LoRA) Hu et al.
+(2022) with rank $r=8$ and scale factor $\alpha=16$.
+We utilize LLMLingua-2 Pan et al.
+(2024) as the token importance metric to generate our compressed CoT training data.
+IRM is evaluated over discrete compression rates in the range $\{0.5,0.6,0.7,0.8,0.9,1.0\}$.
+The $I(k)=w_{a}\cdot e^{\alpha(k-1)}$ was calibrated by fitting to data from controlled error‑injection trials.
+All fine‑tuning used best‑practice hyperparameters.
+We train the models for 5 epochs on both datasets.
+The peak learning rate is set to 5e55e-5, following a cosine decay schedule.
+We use AdamW Loshchilov and Hutter (2017) for optimization, with a warmup ratio of 0.1.
+We implement our training process using the LLaMA-Factory Zheng et al.
+(2024) library.
+During inference, the maximum number of tokens set to 512 for GSM8K and 1024 for MATH.
+A.2. Error‑Injection Experiments
+A.2.1. Error Simulation
+To probe the model’s resilience in a realistic manner, our error-injection methodology does not introduce arbitrary flaws.
+Instead, it is designed to simulate the spontaneous reasoning failures we observed in the model’s own unguided generations, thereby ensuring the ecological validity of our experiment.
+We focused on two distinct classes of these naturally occurring failure patterns.
+The first, Numeric Errors, simulates arithmetic miscalculations.
+The injection process for this error type involves identifying a numerical value in a correct reasoning step and slightly altering it, for instance, by adding or subtracting a small, fixed constant value to mimic a common calculation mistake.
+The second, Symbolic Errors, simulates logical inconsistencies.
+This is achieved by identifying a primary arithmetic operator within a step and systematically replacing it with its mathematical inverse (e.g., + is replaced by -, and * by ÷).
+To maintain procedural integrity, we preemptively implemented constraints to avoid invalid operations, such as division by zero.
+Our overall experimental procedure involved systematically injecting one such controlled error at a specific position k within a baseline of validated, correct reasoning chains.
+Following this injection, the model was prompted to resume the reasoning process from that point, and the accuracy of the final answer was recorded to precisely measure the positional impact of the error.
+A.2.2. Distribution of Results
+Table 5 illustrates the proportion of total steps across all problems.
+We can conclude from the chart that a significant majority of problems are concentrated in the 2-5 step range, with 3-step and 4-step problems being the most prevalent.
+When injecting a numerical error, the accuracy drops are relatively uniform across positions: at step 2 yields a drop of 59.92%, at step 3 a drop of 60.47%, and at step 4 a drop of 61.70%a miscalculation.
+In contrast, symbolic errors induce more severe degradation, and this severity escalates with later injection points.
+Replacement of an operator in the first error step (2/2) reduces the accuracy by 69.67%, while an error in the final step (4/4) inflicts a reduction of 80.80%, with symbolic faults becoming increasingly catastrophic when introduced deeper into the chain.
+Infer Steps Counts Accuracy
+2 786 98.85
+3 2125 96.19
+4 2022 95.75
+5 1180 94.07
+6 484 90.91
+7 137 85.40
+8 58 67.24
+9 7 57.14
+Table 5. Performance analysis by reasoning step count on the GSM8K dataset, illustrating both the frequency of problems per step count and the corresponding model accuracy.
+A.3. Analysis of Late‑Stage Fragility
+Early-stage errors often activate the model’s latent self-correction capability, wherein subsequent reasoning steps can reinterpret and compensate for previous mistakes, leading to a correct final answer.
+This is facilitated by the model maintaining higher semantic entropy and contextual flexibility during the initial stages of reasoning.
+However, as the reasoning chain progresses, the model increasingly commits to a specific semantic trajectory, gradually narrowing its generative space and becoming less sensitive to alternative interpretations.
+We refer to this phenomenon as semantic commitment, a state in which the model’s outputs become rigidly anchored to previously generated content.
+At this point, even minor numerical or symbolic errors near the conclusion are less likely to be detected or corrected, resulting in disproportionately high failure rates.
+Our findings do not aim to refute the cascading failure hypothesis entirely, but rather to refine and contextualize it.
+Previous work suggests that early-stage errors in CoT reasoning tend to propagate forward, compromising subsequent steps and the final answer.
+However, our analysis reveals that the impact of reasoning errors is not uniform.
+We also find that late-stage reasoning steps are disproportionately sensitive to localized perturbations.
+Errors introduced near the conclusion of reasoning are much less likely to be corrected.
+In contrast, early-stage errors may still trigger the model’s latent self-correction mechanisms.
+This indicates that cascading failures do occur, but that late-stage fragility constitutes an additional failure mode not previously emphasized.
+Figure 7. Example generations by ASCoT, using Qwen2.5-Instruct 7B.
+A.4. Generation Examples
+To provide a qualitative illustration of our method’s end-to-end functionality, we present representative examples of reasoning chains after processing by the ASCoT pipeline.
+The Figure 7 displays the corrected output after being processed by ASCoT.
+This transformation from a flawed to a correct reasoning chain highlights the core capabilities of our system.
+It demonstrates the AVM module successfully identifying a high-risk step by analyzing its content quality and positional impact, followed by the MSCE module executing a targeted repair.
+This example visually confirm ASCoT’s ability to structure reasoning but, more importantly, to dynamically detect and rectify internal inconsistencies, ultimately leading to a more trustworthy final answer.
+A.5. Future Work
+We propose several promising future directions.
+A key avenue is the augmentation of MSCE module through integration with external, deterministic oracles like calculators, code interpreters, or knowledge-base APIs.
+By querying these verifiable tools, the correction engine could access ground-truth feedback, significantly enhancing its reliability and moving beyond the constraints of the model’s internal knowledge.
+Another important direction involves developing online adaptive learning mechanisms.
+Such a system would enable ASCoT to continuously self-supervise and refine its performance in real time across various domains, evolving from a static framework into a more dynamic and robust reasoning system.

### v4

@@
-ASCoT: An Adaptive Self-Correction Chain-of-Thought Method for Late-Stage Fragility in LLMs
-Abstract.
-Chain-of-Thought (CoT) prompting has significantly advanced the reasoning capabilities of Large Language Models (LLMs), yet the reliability of these reasoning chains remains a critical challenge.
-A widely held “cascading failure” hypothesis suggests that errors are most detrimental when they occur early in the reasoning process.
-This paper challenges that assumption through systematic error-injection experiments, revealing a counter-intuitive phenomenon we term “Late-Stage Fragility”: errors introduced in the later stages of a CoT chain are significantly more likely to corrupt the final answer than identical errors made at the beginning.
-To address this specific vulnerability, we introduce the Adaptive Self-Correction Chain-of-Thought (ASCoT) method.
-ASCoT employs a modular pipeline in which an Adaptive Verification Manager (AVM) operates first, followed by the Multi-Perspective Self-Correction Engine (MSCE).
-The AVM leverages a Positional Impact Score function that assigns different weights based on the position within the reasoning chains, addressing the Late-Stage Fragility issue by identifying and prioritizing high-risk, late-stage steps.
-Once these critical steps are identified, the MSCE applies robust, dual-path correction specifically to the failure parts.
-Extensive experiments on benchmarks demonstrate our superiority over previous methods.
-ASCoT reduces token usage by 21%/30% for the LLaMA-3.1-8B on GSM8K/MATH-500 benchmarks, with minimal accuracy drops of only 1.1% and 1.8%, achieving new sota performance.
-Key words and phrases:
-Large Language Models, Common-Sense Reasoning, Interpretability and Evaluation of NLP Models, Prompt Engineering
-1. Introduction
-Figure 1.
-The top part (a) illustrates the standard CoT reasoning process, which often generates redundant outputs and lacks an effective self-correction mechanism when errors occur.
+Not All Errors Are Created Equal: ASCoT Addresses Late-Stage Fragility in Efficient LLM Reasoning
+Abstract
+While Chain-of-Thought (CoT) prompting empowers Large Language Models (LLMs), ensuring reasoning reliability remains an open challenge.
+Contrary to the prevailing cascading failure hypothesis which posits that early errors are most detrimental, we identify a counter-intuitive phenomenon termed Late-Stage Fragility: errors introduced in later reasoning stages are significantly more prone to corrupting final answers.
+To address this, we introduce ASCoT (Adaptive Self-Correction Chain-of-Thought), a method harmonizing efficiency with robust verification.
+ASCoT first employs semantic pruning to compress redundant steps, then utilizes an Adaptive Verification Manager (AVM) to prioritize high risk, late-stage steps via a positional impact score, triggering a Multi-Perspective Self-Correction Engine (MSCE) only when necessary.
+Experiments on GSM8K and MATH-500 demonstrate that ASCoT effectively reallocates computational resources: it reduces token usage by 21%–30% for LLaMA-3.1-8B with negligible accuracy drops ($1.8\%$), achieving a superior trade-off between inference efficiency and reasoning fidelity.
+Not All Errors Are Created Equal: ASCoT Addresses Late-Stage Fragility in Efficient LLM Reasoning
+[…]
+1 Introduction
+Figure 1: The top part (a) illustrates the standard CoT reasoning process, which often generates redundant outputs and lacks an effective self-correction mechanism when errors occur.
 The bottom part (b) depicts the ASCoT method, which incorporates a robust mechanism for self-correction, addressing potential reasoning flaws at each stage.
-The advent of Chain-of-Thought (CoT) has endowed Large Language Models (LLMs) with remarkable multi-step reasoning capabilities Kojima et al.
-(2022), enabling them to deconstruct complex problems into a sequence of manageable sub-problems Yao et al.
-(2023); Shinn et al.
+The advent of CoT has endowed LLMs with remarkable multi-step reasoning capabilities Kojima et al.
+(2022), enabling them to deconstruct complex problems into a sequence of manageable sub-problems Shinn et al.
 (2023).
@@
 However, the success of CoT and its variants has introduced challenges: prohibitive computational costs from long reasoning chains Chen et al.
-(2025) and the inherent brittleness of the reasoning process itself Zhang et al.
-(2024); Ke et al.
+(2025) and the inherent brittleness of the reasoning process itself Ke et al.
 (2025).
-A single flaw is often sufficient to invalidate an entire process of thought, raising significant concerns about the trustworthiness of these models Chowdhury and Caragea (2025).
-A widely held belief within the research community, often termed the “cascading failure” hypothesis, posits that the primary threat to CoT integrity lies in early-stage errors Tyen et al.
+A single flaw is often sufficient to invalidate an entire process of thought, raising significant concerns about the trustworthiness of these models Pan et al.
+(2024).
+A widely held belief within the research community, often termed the cascading failure hypothesis, posits that the primary threat to CoT integrity lies in early-stage errors Tyen et al.
 (2023).
@@
 We discover that errors introduced in later stages are more likely to lead to an incorrect final answer.
-We term this phenomenon “Late-Stage Fragility”.
+We term this phenomenon Late-Stage Fragility.
 An early error often appears to trigger the model’s latent self-correction mechanisms, allowing it to recover and still reach the correct solution.
-As the model progresses steadily, its reasoning process gains a form of “semantic commitment”, reducing relevant information entropy.
+As the model progresses steadily, its reasoning process gains a form of semantic commitment, reducing relevant information entropy.
 At the later stages, the model is unable to evaluate the validity of the final calculations.
 This empirical discovery suggests that a significant portion of efforts to secure CoT reasoning may be misdirected.
-As illustrated in Figure 1(a), a standard CoT process can be both inefficient due to redundancy and unreliable due to its inability to correct these late-stage errors.
+As illustrated in Figure 1, a standard CoT process can be both inefficient due to redundancy and unreliable due to its inability to correct these late-stage errors.
 Building on a rethinking of how we evaluate and how we correct LLM reasoning, we introduce the ASCoT method.
-In contrast to standard methods, ASCoT is designed to address both challenges simultaneously.
-As shown in Figure 1(b), it first employs a semantic pruning stage to enhance efficiency, followed by a targeted, risk-aware correction mechanism that focuses computational resources precisely on the most vulnerable steps to ensure a correct final answer.
-To sum up, our contributions are summarized as follows:
-  • As far as we know, this is the first to identify and quantify the “Late-Stage Fragility” phenomenon in CoT reasoning, demonstrating that errors in later steps are significantly more detrimental than those in earlier steps.
-  • We introduce ASCoT, a novel method specifically designed to counteract this fragility by precisely identifying and robustly correcting the high-impact, late-stage errors that are most critical to reasoning accuracy.
-  • Our extensive experiments validate the effectiveness of ASCoT.
-We demonstrate that ASCoT achieves outstanding accuracy on the GSM8k and MATH benchmarks, validating our approach of diagnosing the specific failure mode before engineering a targeted solution.
-2. Preliminary
-This section establishes the foundational concepts that motivate our work, offering a review of some effective methods for identifying token importance within CoT.
+In contrast to standard methods, ASCoT is designed to address both challenges simultaneously, it first employs a semantic pruning stage to enhance efficiency, followed by a targeted, risk-aware correction mechanism that focuses computational resources precisely on the most vulnerable steps to ensure a correct final answer.
+Our contributions are summarized as follows:
+  • To the best of our knowledge, this is the first to identify and quantify the Late-Stage Fragility phenomenon in CoT reasoning, demonstrating that errors in later steps are significantly more detrimental than those in earlier steps.
+  • We introduce ASCoT, a novel method specifically designed to counteract this fragility by precisely identifying and robustly correcting the high-impact, late-stage errors.
+  • Our extensive experiments validate the effectiveness of ASCoT. We demonstrate that ASCoT achieves outstanding accuracy on the GSM8k and MATH benchmarks.
+2 Related Work
+2.1 LLM Reasoning & Efficient Thinking
+CoT prompting significantly advances LLM reasoning by decomposing complex problems into intermediate steps Liu et al.
+(2024b); Team (2024).
+To mitigate the associated computational costs, recent research explores diverse efficiency-focused strategies, including prompt-based constraints ($e.g.,$ token budgets) Xu et al.
+(2025); Han et al.
+(2024), inference-time dynamic early-exit mechanisms Ding et al.
+(2025), and CoT distillation Wadhwa et al.
+(2024); Zhang et al.
+(2026).
+2.2 CoT Compression and Pruning
+The inherent verbosity of CoT incurs significant computational overhead Wingate et al.
+(2022) and can induce the overthinking phenomenon, where excessive reasoning introduces errors Kumar et al.
+(2025).
+To mitigate this, recent research focuses on compression strategies Xuefei et al.
+(2023).
+Approaches vary from token-level pruning based on semantic importance Zhu et al.
+(2025); Yang et al.
+(2025) to structural methods like Prune-on-Logic Liu et al.
+(2025), which filters steps via logic graphs.
+Notably, even heuristic truncation Wang et al.
+(2024) has demonstrated efficacy, suggesting that exhaustive reasoning paths are often redundant.
+2.3 Self-correction Mechanism
+Self-correction mechanisms enhance LLM reliability by iterating through generation, feedback, and refinement phases Kamoi et al.
+(2024).
+Current approaches bifurcate based on the feedback source.
+Intrinsic self-correction relies on the model’s internal critique capabilities without external tools Yang et al.
+(2025); Zhu et al.
+(2025).
+However, these methods often struggle to identify errors due to the scarcity of correction examples in training data Liu et al.
+(2024a).
+In contrast, extrinsic self-correction leverages external signals, such as verifier models Kim et al.
+(2025) or consistency checks, to rigorously assess step-wise correctness.
+3 Preliminary
+This section establishes the foundational concepts that motivate our work, offering a review of some effective methods for identifying token importance.
 Efficient CoT in LLM Reasoning A critical question for CoT efficiency is whether every token in different position contributes equally to the final answer Xia et al.
@@
 Several early approaches Hou et al.
-(2022); Zhang et al.
-(2023); Chen et al.
-(2024b) measure importance based on the self-information or surprisal from a unidirectional language model:
+(2022) measure importance based on the self-information or surprisal from a unidirectional language model:
 $I_{1}(x_{i})=-\log P(x_{i}|x_{<i};\theta_{\mathcal{M}_{L}}),\quad$ (1)
@@
 To address this, more advanced methods like LLMLingua-2 Pan et al.
-(2024) utilize a bidirectional language model Jawahar et al.
-(2019); Devlin et al.
+(2024) utilize a bidirectional language model Devlin et al.
 (2019) ($e.g.,$ BERT-style) to reframe the task as token classification.
-In this paradigm, the importance of a token $x_{i}$ is defined as the probability that a trained classifier assigns it an ”essential” label ($y_{i}=1$), conditioned on the entire sequence $x$:
+In this paradigm, the importance of a token $x_{i}$ is defined as the probability that a trained classifier assigns it an essential label ($y_{i}=1$), conditioned on the entire sequence $x$:
 $I_{2}(x_{i})=P(y_{i}=1|x;\theta_{\mathcal{M}{cls}}),$ (2)
@@
 For example, tokens comprising mathematical equations or key entities Ma et al.
-(2025) are identified as critical, whereas semantic connectors such as so” and since” contribute less to the final outcome of the reasoning Hou et al.
+(2025) are identified as critical, whereas semantic connectors such as “so” and “since” contribute less to the final outcome of the reasoning  Hou et al.
 (2025).
@@
 A single error can invalidate an entire reasoning process Guo et al.
-(2017); Tyen et al.
 (2017).
-The conventional understanding of this problem is rooted in the “cascading failure” hypothesis, which posits that errors are most damaging when they occur early in the reasoning chain Jiang et al.
+The conventional understanding of this problem is rooted in the cascading failure hypothesis, which posits that errors are most damaging when they occur early in the reasoning chain Jiang et al.
 (2025).
@@
 In this study, through systematic error-injection experiments, we observe that errors introduced in the later stages of a reasoning chain are significantly more detrimental to the final outcome than identical errors introduced at the beginning.
-Figure 2.
-Illustration of ASCoT Pipeline.
+Figure 2: Illustration of ASCoT Pipeline.
 ASCoT generates the $CoT_{initial}$ from the target LLM, which is then compressed by IRM module to a predefined ratio $\gamma$.
@@
 Finally, ASCoT enables reasoning on new problems with the specified compression ratio $\gamma$.
-Algorithm 1 ASCoT Mechanism
-Input: Initial chain-of-thought (CoT) $CoT_{initial}$
-Output: Corrected $CoT_{final}$
-[…]
-[…]
-3:for each step $t_{k}$ in $CoT_{pruned}$ do
-4:  $is\_flagged\leftarrow\mathrm{AVM}(context,t_{k})$
-5:  if is_flaggedis\_flagged then
-6:   $t_{k}^{\mathrm{corr}}\leftarrow\mathrm{MSCE}(context,t_{k})$
-7:   append $t_{k}^{\mathrm{corr}}$ to $CoT_{final}$
-8:   $context\leftarrow context\,\|\,t_{k}^{\mathrm{corr}}$
-9:  else
-10:   append $t_{k}$ to $CoT_{final}$
-11:   $context\leftarrow context\,\|\,t_{k}$
-12:  end if
-13:end for
-14:return $CoT_{final}$
-3. Method
+4 Method
 In this section, we detail the architecture of our method.
 ASCoT generates the $CoT_{initial}$ based on the target LLM.
-The $CoT_{initial}$ is then processed through the IRM module, reducing the output to a fixed ratio $\gamma$.
-The remaining CoT is passed through the AVM module, which performs a rigorous verification process.
-If the model’s risk score for any CoT step exceeds a predefined threshold $\tau$, indicating a potential error, the corresponding CoT step is forwarded to the MSCE module.
+The $CoT_{initial}$ is then processed through the Intelligent Routing Mechanism (IRM) module, reducing the output to a fixed ratio $\gamma$.
+The remaining CoT is passed through the Adaptive Verification Manager (AVM) module, which performs a rigorous verification process.
+If the model’s risk score for any CoT step exceeds a predefined threshold $\tau$, indicating a potential error, the corresponding CoT step is forwarded to the Multi-Perspective Self-Correction Engine (MSCE) module.
 The MSCE module applies internal and external validation techniques to correct errors, resulting in the corrected $CoT_{final}$.
@@
 We illustrate the training and inference process in Figure 2, respectively.
-We also provide Pseudocode 1 for ASCoT.
-3.1. Intelligent Routing Mechanism
+We also provide Algorithm 1 for ASCoT.
+4.1 Intelligent Routing Mechanism
 The IRM serves as the initial processing stage, responsible for preliminary CoT pruning and compression.
@@
 Based on a predefined compression ratio $\gamma$, tokens with scores below the corresponding threshold are pruned, resulting in a compressed CoT, namely the final answer.
-3.2. Adaptive Verification Manager
+4.2 Adaptive Verification Manager
 AVM dynamically identifies which steps are most likely to be incorrect and then computes a Risk Score ${R}(t_{k})$ for each step $t_{k}$ by integrating two key assessments: the step’s confidence assessment $Q(t_{k})$ and its positional impact $I(k)$.
@@
 Figure 3 illustrates this evaluation process, where an initial CoT with a calculation error is analyzed.
-Figure 3.
-The details of verification and self-correction.
-AVM and MSCE working together to correct a faulty step $t_{k}$.
-AVM computes $R(t_{k})$.
-MSCE then applies intrinsic and extrinsic correction to produce the final corrected $CoT_{final}$.
-Positional Impact Score To quantitatively model the “Late-Stage Fragility” phenomenon observed in our error-injection experiments (Section 4.3), we introduce the Position Impact Function $I(k)$.
+Positional Impact Score To quantitatively model the Late-Stage Fragility phenomenon observed in our error-injection experiments (Section 4.3), we introduce the Position Impact Function $I(k)$.
 This function is an empirical model fitted to our experimental data, designed to capture the escalating impact of errors at later reasoning stages.
@@
 The function $I(k)$ is then fitted to these empirical data points.
-Details of the analysis are available in Appendix.
+Details of the analysis are available in Appendix E.
 Finally, AVM combines these factors into a single Risk Score:
 ${R}(t_{k})=I(k)\,\times\bigl(1-Q(t_{k})\bigr).$ (9)
-This score is then compared against a tunable system threshold $\tau$. If $R(t_{k})>\tau$, the step is flagged as high-risk and passed to the MSCE for correction.
-Table 1.
-Experimental results of ASCoT on LLaMA-3.1-8B-Instruct.
-We report Accuracy, average CoT token count (Tokens) and actual compression ratio (ActRatio) for comparison.
-The compression ratio $\gamma$ is set to 0.51.00.5-1.0, respectively.
-Methods Ratio GSM8K MATH-500
-Accuracy $\uparrow$ Tokens $\downarrow$ ActRatio Accuracy $\uparrow$ Tokens $\downarrow$ ActRatio
-Original 1.0 86.2(0.0↓) 213.17 1.00 48.6(0.0↓) 502.60 1.00
-Prompt 0.9 84.1(2.1↓) 226.37 1.06 48.6(0.0↓) 468.04 0.93
-0.7 84.9(1.3↓) 209.39 0.98 48.4(0.4↓) 472.13 0.94
-0.5 83.7(2.5↓) 188.82 0.89 47.8(0.4↓) 471.11 0.94
-Truncation 0.9 70.2(26.0↓) 202.06 0.95 47.8(0.8↓) 440.33 0.88
-0.7 25.9(60.3↓) 149.97 0.70 45.0(3.6↓) 386.89 0.77
-0.5 7.0(79.2↓) 103.69 0.49 27.4(21.2↓) 283.70 0.56
-ASCoT 1.0 86.9(0.7↑) 214.24 1.00 48.8(0.2↑) 478.80 0.95
-0.9 86.1(0.1↓) 197.65 0.92 47.6(1.0↓) 427.15 0.84
-0.8 85.1(1.1↓) 169.96 0.79 47.2(1.4↓) 376.72 0.74
-0.7 83.7(2.5↓) 151.32 0.71 46.8(1.8↓) 334.26 0.67
-0.6 81.6(4.6↓) 128.38 0.60 41.8(6.8↓) 309.69 0.61
-0.5 79.5(6.7↓) 118.57 0.55 40.1(8.5↓) 288.94 0.57
-3.3. Multi-Perspective Self-Correction Engine
+This score is then compared against a tunable system threshold $\tau$. If $R(t_{k})>\tau$, the step is flagged as high risk and passed to the MSCE.
+4.3 Multi-Perspective Self-Correction Engine
 Once a step $t_{k}$ is flagged by the AVM, the MSCE is activated to perform a robust and efficient correction.
-To counteract a model struggles to correct its own mistakes when using a single line of reasoning, the MSCE employs a dual-path correction strategy: Intrinsic Correction (Self-Reflection), the model is presented with the context up to step $s_{(k-1)}$ and its own flawed step $t_{k}$, and is explicitly prompted to review and amend it.
-Extrinsic Correction (Independent Generation), the model is given the context up to $s_{(k-1)}$ but without the flawed step $t_{k}$.
+To counteract a model struggles to correct its own mistakes when using a single line of reasoning, the MSCE employs a dual-path correction strategy: Intrinsic Correction, the model is presented with the context up to step $s_{(k-1)}$ and its own flawed step $t_{k}$, and is explicitly prompted to review and amend it.
+Extrinsic Correction, the model is given the context up to $s_{(k-1)}$ but without the flawed step $t_{k}$.
 Upon generating two candidate corrections, the MSCE employs a selection mechanism to choose the superior one.
@@
 The candidate with the higher quality score is then integrated into the reasoning chain.
-4. Experiment
-4.1. Experimental Setup
-Models and Datasets We primarily evaluate our method using the LLaMA-3.1-8B-Instruct Ge et al.
+Figure 3: The details of verification and self-correction.
+AVM and MSCE working together to correct a faulty step $t_{k}$.
+AVM computes $R(t_{k})$.
+MSCE then applies intrinsic and extrinsic correction to produce the final corrected $CoT_{final}$.
+Table 1: Experimental results of ASCoT on LLaMA-3.1-8B-Instruct.
+We report Accuracy, average CoT token count (Tokens) and actual compression ratio (ActRatio) for comparison.
+The compression ratio $\gamma$ is set to 0.51.00.5-1.0.
+Methods Ratio GSM8K MATH-500
+Accuracy $\uparrow$ Tokens $\downarrow$ ActRatio Accuracy $\uparrow$ Tokens $\downarrow$ ActRatio
+Original 1.0 86.2($0\downarrow$) 213.17 1.00 48.6($0\downarrow$) 502.60 1.00
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+5 Experiment
+5.1 Experimental Setup
+Models and Datasets We primarily evaluate our method using the LLaMA-3.1-8B-Instruct Dubey et al.
 (2024) and Qwen2.5-Instruct series (3B, 7B, 14B) Team (2024)models.
@@
 (2022) for all fine-tuning procedures, setting the rank $r=8$ and scaling parameter $\alpha=16$ to ensure efficient training.
-Additional implementation details and configurations can be found in the Appendix.
+Additional implementation details and configurations can be found in the Appendix D.
 Evaluation Metrics We evaluate the performance of ASCoT using three primary metrics: Accuracy (%), which represents the percentage of problems for which the final answer is correct.
 Tokens, the average number of tokens generated in the reasoning chain, measuring computational cost and efficiency during the reasoning process.
-ActRatio, the actual compression ratio of ASCoT, reflecting the real efficiency of the token compression process.
-4.2. Main Results
-We evaluate the end-to-end performance of ASCoT against two prevalent CoT length control techniques.
-The first is Prompt, appends an explicit instruction (e.g., “Please reduce your CoT output by 50%”) to the model’s input, thus guiding the LLM to generate a shorter reasoning trace.
-The second termed Truncation, imposes a fixed upper bound on the number of output tokens during inference.
-The results of LLaMA-3.1-8B-Instruct are summarized in Table 1.
-As shown, prompt‐based reduction consistently fails to meet the target ratio: even at a nominal 0.5 budget it yields an actual compression ratio of 0.94 on MATH-500, with only marginal savings in CoT tokens.
-Truncation strictly enforces the specified ratio but at the cost of dramatic performance degradation: at a budget of 0.5, the accuracy on GSM8K plummets by 79.2 points and by 21.2 points on MATH-500, rendering it unsuitable for tasks requiring robust multistep reasoning.
-In contrast, ASCoT closely matches the specified budget (e.g., ActRatio = 0.79 at Ratio = 0.8 on GSM8K and ActRatio = 0.61 at Ratio = 0.6 on MATH-500) while maintaining robust reasoning performance.
-In a full budget (Ratio = 1.0), ASCoT marginally outperforms the baseline, achieving 86.9% on GSM8K (+0.7) and 48.8% on MATH-500 (+0.2).
-In at a 0.5 budget, it maintains respectable accuracies of 79.5% (–6.7) on GSM8K and 40.1% (–8.5) on MATH-500, respectively.
-These results demonstrate that ASCoT strikes a favorable balance between compression and accuracy, validating its effectiveness.
-Figure 4.
-Performance of ASCoT on the Qwen2.5-Instruct Series.
+ActRatio, the actual compression ratio of ASCoT, reflecting the real efficiency of the token compression.
+Figure 4: Performance of ASCoT on the Qwen2.5-Instruct Series.
 The results for the 3B, 7B, and 14B models are shown under various compression ratios and compared against the original, uncompressed baseline.
-Figure 5. Performance comparison of ASCoT with varying maximum length constraints. We adjust the maximum length budget when evaluating our method on MATH-500.
-We also provide the results of Qwen2.5‐Instruct series models as Table 2.
-Across the three model scales (3B, 7B, and 14B), ASCoT consistently maintains original performance while achieving substantial CoT compression.
-For the 3B model, ASCoT at 0.9 compression incurs only a 0.5 point drop in accuracy, while reducing token usage by 16% (314.87 $\rightarrow$ 262.31).
-As the compression ratio decreases to 0.5, the accuracy degrades from 83.5% to 74.9%, but the token count falls sharply to 170.87(reducing 45.7%).
-Similar trends hold for the 7B and 14B variants: at Ratio = 0.9, both scales suffer less than a 0.5 point accuracy drop (for the 14B model, there is even a slight improvement in performance), yet token counts shrink by about 15%.
-Even under aggressive compression (ratio = 0.5), the 14B model retains more than 91.6% accuracy, representing only a 1.5 point reduction, while halving the original token budget.
-The detailed experimental results can be found in the appendix.
-Table 2. Experimental results of the Qwen2.5-Instruct series evaluated on GSM8K. We also provide Accuracy, Tokens and ActRatio.
+5.2 Main Results
+We benchmark ASCoT against two standard length-control baselines: instruction-based Prompting and hard Truncation. The comparative results on LLaMA-3.1-8B are presented in Table 1.
+Comparison with Baselines Existing methods struggle to balance compression controllability with reasoning fidelity.
+Prompting fails to strictly adhere to token budgets.
+Truncation also incurs catastrophic performance degradation when it enforces strict length constraints.
+In contrast, ASCoT demonstrates precise controllability and robust performance.
+It aligns closely with target ratios (e.g., ActRatio 0.79 at a 0.8 target) while preserving reasoning accuracy.
+Notably, at a full budget ($\gamma=1.0$), ASCoT marginally outperforms the original baseline (+0.7% on GSM8K), suggesting that our verification mechanism effectively filters noise.
+Even under aggressive compression ($\gamma=0.5$), ASCoT maintains a respectable 79.5% accuracy on GSM8K, significantly outperforming truncation based methods.
+Scalability Analysis Table 2 details the performance across the Qwen2.5-Instruct series (3B, 7B, 14B) to evaluate scalability.
+ASCoT achieves substantial token reduction with minimal accuracy loss across all scales.
+For the 3B model, a 0.9 compression ratio reduces token usage by approximately 16% with only a 0.5 point accuracy drop.
+This resilience becomes more pronounced as model size increases.
+The 14B model exhibits exceptional robustness.
+Even when halving the token budget, it retains 91.6% accuracy (a mere 1.5% drop).
+This trend suggests a positive correlation between model scale and reasoning redundancy, larger models tend to generate more steps that ASCoT effectively identifies and prunes without compromising the final answer, demonstrating its potential to significantly lower inference costs for large-scale deployments without sacrificing reliability.
+Figure 5: Performance comparison of ASCoT with varying maximum length constraints. We adjust the maximum length budget when evaluating our method on MATH-500.
+Figure 4 illustrates ASCoT’s robust scalability across the Qwen2.5-Instruct series.
+While the 3B model incurs a negligible 0.5-point accuracy drop at 17% compression, this resilience amplifies with model size.
+The 7B variant maintains over 86% accuracy even with a 50% token reduction.
+Most notably, the 14B model demonstrates exceptional efficiency retention: halving the token budget results in a mere 1.5% accuracy decline.
+This trend indicates a positive correlation between model capacity and reasoning redundancy, suggesting that ASCoT effectively exploits the richer semantic structures inherent in larger models to maximize pruning without information loss.
+This confirm ASCoT’s capability to balance computational efficiency and reasoning fidelity across varying model scales.
+5.3 Error Propagation of Controlled Injection
+To interrogate the cascading failure hypothesis, we conducted a controlled error-injection study on GSM8K using Qwen2.5-7B.
+We injected two classes of perturbations mimicking naturally occurring mistakes: Numeric (value deviation) and Symbolic (operator inversion).
+Results in Table 4 reveal a Late-Stage Fragility.
+Contrary to the expectation that early errors propagate most severely, we observe that impact escalates with position.
+For numeric errors, a miscalculation at the final step (4/4) causes a catastrophic 51.69% accuracy drop, significantly outpacing early-stage errors (14.64% at Step 2).
+Symbolic errors amplify this trend, with late-stage injections (Step 3/3) causing the most profound degradation (56.78%).
+These findings suggest that models become increasingly committed to their reasoning trajectory, losing the capacity to self-correct as they approach the solution.
+Distribution details of the inference steps are provided in Table 3.
+Table 2: Experimental results of the Qwen2.5-Instruct series evaluated on GSM8K. We report Accuracy (%), Tokens and ActRatio.
 Scale Methods Ratio Accuracy Tokens ActRatio
-3B Original 83.7(0.0↓) 314.87
-ASCoT 1.0 83.5(0.2↓) 318.32 1.01
-0.9 83.2(0.5↓) 262.31 0.83
-0.8 82.0(1.7↓) 250.75 0.79
-0.7 80.1(3.6↓) 233.15 0.74
-0.6 77.6(6.1↓) 199.44 0.63
-0.5 74.9(8.8↓) 170.87 0.54
-7B Original 91.4(0.0↓) 297.83
-ASCoT 1.0 91.8(0.4↑) 294.74 0.99
-0.9 91.3(0.1↓) 251.71 0.85
-0.8 90.4(1.0↓) 234.95 0.79
-0.7 90.3(1.1↓) 216.73 0.73
-0.6 88.2(3.2↓) 178.28 0.60
-0.5 86.3(5.1↓) 150.98 0.51
-14B Original 93.1(0.0↓) 313.11
-ASCoT 1.0 93.3(0.2↑) 312.30 0.99
-0.9 93.6(0.5↑) 269.76 0.86
-0.8 93.5(0.4↑) 246.57 0.79
-0.7 93.7(0.6↑) 221.17 0.71
-0.6 92.8(0.3↓) 177.40 0.57
-0.5 91.6(1.5↓) 158.21 0.51
-Figure 4 demonstrates that ASCoT consistently achieves substantial CoT compression across all tested Qwen2.5-Instruct series models (3B, 7B, and 14B).
-For the 3B model, a moderate compression that reduces token usage by nearly 17% (from 314.87 to 262.31 tokens) incurs only a negligible 0.5-point drop in accuracy.
-This trend is further amplified in the larger models.
-With the 7B variant, a 15% token reduction results in a minimal 0.1-point accuracy drop, and a nearly 50% reduction in tokens still maintains accuracy above 86%.
-Notably, the 14B model showcases ASCoT’s exceptional ability to enhance efficiency.
-Even when halving the original token budget (a 50% reduction to 158.21 tokens), the accuracy drop is a mere 1.5 percentage points.
-These results validate ASCoT’s effectiveness in striking a favorable balance between computational efficiency and reasoning fidelity, demonstrating its robustness and scalability across different model sizes.
-4.3. Error Propagation via Controlled Injection
-To quantitatively investigate the dynamics of error propagation in CoT reasoning, we conducted a comprehensive and systematic error-injection study to rigorously test the “cascading failure” hypothesis.
-Our experimental setup begins with a baseline of correct CoT solutions for problems from the GSM8K benchmark.
-For each correct reasoning chain, we first introduce two distinct classes of perturbations into the reasoning steps of the correct CoT sequences.
-Then, we assess and evaluate how these perturbations affect the overall reasoning process.
-The injected errors are not manually added incorrect tokens, but simulated based on real symbolic and numerical mistakes observed during the model’s own generation.
-By replicating these naturally occurring failure patterns in a controlled manner, we enable systematic analysis of how the position and type of such errors affects the final reasoning outcome.
-Numeric errors were introduced to simulate arithmetic miscalculations by slightly altering a single number in the reasoning process (e.g., adding or subtracting a fixed constant value).
-Symbolic errors were introduced to simulate logical inconsistencies by replacing a primary arithmetic operator with its inverse (e.g.
-+ to -, $\times$ to $\div$).
-We preemptively set strict constraints to avoid division by zero.
-Table 4 reports the impact of single‐step perturbations generated by Qwen2.5-Instruct 7B.
-For numeric errors, the accuracy drops gradually across positions, with a miscalculation at step 2 causing a decrease of 14.64%, at step 3 a 29.56% drop, and at step 4 a significant reduction of 51.69%.
-In contrast, symbolic errors induce more severe performance degradation, especially when injected later in the reasoning process.
-A symbolic error at step 2 results in a large drop of 43.97%,.
-However, injecting a symbolic error at step 3 (3/3) results in an accuracy reduction of 56.78%.
-These results emphasize the susceptibility of CoT reasoning to errors, particularly with symbolic mistakes, which cause a more pronounced decline in performance as the reasoning progresses.
-Table 3 reports initial accuracy before error injection.
-All the above data are from dataset GSM8K.
-Table 3. Accuracy for Different Inference Steps in the GSM8K Dataset, using Qwen2.5-Instruct 7B.
+3B Original 83.7($0\downarrow$) 314.87
+[…]
+[…]
+[…]
+[…]
+[…]
+0.5 74.9($8\downarrow$) 170.87 0.54
+7B Original 91.4($0\downarrow$) 297.83
+[…]
+[…]
+[…]
+0.7 90.3($1\downarrow$) 216.73 0.73
+[…]
+[…]
+14B Original 93.1($0\downarrow$) 313.11
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+5.4 Analysis
+Length Budget Analysis To further investigate ASCoT’s efficiency, we analyze its performance under a fixed token budget.
+As illustrated in Figure 5, we assess our approach under two distinct length-budget settings.
+In the first, the maximum generation length is scaled by the compression ratio (ranging from 0.5 to 1.0).
+In the second, we retain the original LLM budget ($e.g.,$ 512 tokens in GSM8K and 1024 tokens in MATH) regardless of compression.
+The results reveal that under the unchanged budget, our method consistently surpasses the baseline: accuracy rises from 42.6% to a peak of 49.4%.
+Notably, at compression ratios of 0.8 and 0.9, we observe absolute accuracy gains of 0.8 and 0.6 percentage points over the baseline in a smaller tokens, underscoring the potential of compressed CoT reasoning within the same budget.
+The Late-Stage Fragility of CoT Our findings refine the traditional cascading failure model by uncovering Late-Stage Fragility phenomenon.
+While early-stage errors often trigger latent self-correction mechanisms due to high contextual entropy, identical errors introduced in the final steps are significantly more prone to persist uncorrected (qualitative illustration provided in Appendix C).
+This suggests that the model forms a rigid semantic commitment as reasoning progresses.
+Consequently, while cascading failures undeniably exist, Late-Stage Fragility constitutes a distinct, critical failure mode where the model lacks the flexibility to revisit and rectify terminal errors.
+Case Study To provide a clear and detailed qualitative understanding of our method, we present a case study.
+Figure 3 illustrates a qualitative analysis where a standard model correctly structures the reasoning but commits a critical arithmetic error by calculating $125\times 8$ as $1005$.
+This flaw propagates to an incorrect conclusion.
+ASCoT effectively addresses this issue as the AVM module flags the step due to low Logical Validity while acknowledging its structural soundness.
+Crucially, this granular diagnosis prevents the unnecessary discarding of the entire reasoning chain.
+Instead, the MSCE executes a targeted dual-path repair to synthesize a corrected candidate, rectifying the calculation to $1000$, thereby enabling the model to recover and deduce the correct total of $1015$.
+Table 3: Accuracy for Different Inference Steps in the GSM8K Dataset, using Qwen2.5-Instruct 7B.
 Steps Correct Total Accuracy
@@
 8 39 58 67.24
-9 4 7 57.14
-10 11 11 100.00
-11 2 2 100.00
-4.4. Analysis
-Length Budget Analysis To further investigate ASCoT’s efficiency, we analyze its performance under a fixed token budget.
-As illustrated in Figure 5, we assess our approach under two distinct length-budget settings.
-In the first, the maximum generation length is scaled by the compression ratio (ranging from 0.5 to 1.0).
-In the second, we retain the original LLM budget (e.g., 512 tokens in GSM8K and 1024 tokens in MATH) regardless of compression.
-The results reveal that under the unchanged budget, our method consistently surpasses the baseline: accuracy rises from 42.6% to a peak of 49.4%.
-Notably, at compression ratios of 0.8 and 0.9, we observe absolute accuracy gains of 0.8 and 0.6 percentage points over the baseline in a smaller tokens, underscoring the potential of compressed CoT reasoning within the same budgetary constraints.
-The Late-Stage Fragility of CoT Our findings reveal a phenomenon that challenges the traditional cascading failure model.
-To illustrate this conceptually, Figure 6 contrasts the model’s ability to recover from an identical type of error when introduced at different stages of the reasoning process.
-An early-stage error, while potentially disruptive, was often corrected or mitigated by the model’s subsequent reasoning steps.
-However, an identical error introduced in the final or penultimate step was significantly more likely to go uncorrected, directly leading to a wrong answer.
-Our findings do not aim to refute the cascading failure hypothesis entirely, but rather to refine and contextualize it.
-Previous work suggests that early-stage errors in CoT reasoning tend to propagate forward, compromising subsequent steps and the final answer.
-However, our analysis reveals that the impact of reasoning errors is not uniform.
-Errors introduced near the conclusion of reasoning are notably much less likely to be corrected.
-This indicates that cascading failures do occur, but that late-stage fragility constitutes an additional failure mode not previously emphasized.
-Figure 6.
-A illustration of “Late-Stage Fragility”.
-The model recovers from an early-stage calculation error, but fails when an identical type of error is introduced in the final step.
-The tokens that occur related errors are highlighted in orange.
-Case Study To provide a clear and detailed qualitative understanding of our method, we present a case study, illustrated in Figure 3.
-The problem asks for the total number of apples Bob picked over nine days.
-As shown in the Initial CoT panel, a standard model correctly structures the reasoning but commits a calculation error in Step 2, incorrectly stating that 125 * 8 = 1,005.
-This single subtle yet critical flaw propagates directly to the final step, resulting in an incorrect answer of 1,020.
-The Evaluation panel demonstrates the intervention of AVM module.
-The AVM performs a multi-dimensional assessment of the flawed step.
-It correctly identifies the lack of Logical Validity (0/1) due to the arithmetic error.
-Simultaneously, it recognizes that the step’s Factual Groundedness and Semantic Clarity & Process Utility are sound.
-This evaluation allows the AVM to pinpoint the exact failure without discarding the entire step’s valid structure.
-Because the step’s risk score exceeds the threshold, the MSCE is activated.
-As shown in the Final result panel, the MSCE provides the corrected calculation (125 * 8 = 1,000) using a dual-path correction mechanism.
-This targeted repair allows the model to obtain the correct final answer of 1,015.
-Figure 3 illustrates this evaluation process, where an initial CoT with a calculation error is analyzed.
-5. Related Work
-5.1. LLM Reasoning & Efficient Thinking
-The advent of CoT prompting marked a significant leap in the reasoning capabilities of LLMs Marconato et al.
-(2023); Ding et al.
-(2024); Liu et al.
-(2024a).
-CoT decomposes complex problems into a sequence of manageable intermediate steps, dramatically improving performance on arithmetic, commonsense, and symbolic reasoning tasks Team (2024).
-These efficiency-focused methods can be broadly categorized: Prompt-based strategies Xu et al.
-(2025) aim to guide the model toward more concise reasoning through specific instructions in the prompt itself, such as setting a token budget Han et al.
-(2024).
-Inference-time dynamic reasoning Ding et al.
-(2025) includes early-exit mechanisms that terminate the reasoning process once a certain confidence threshold is met, avoiding unnecessary computation.
-CoT Distillation Wadhwa et al.
-(2024); Wang et al.
-(2023) focuses on transferring the reasoning ability of a large “teacher” model to a smaller “student” model.
-Table 4.
-Error-injection results on GSM8K, using Qwen2.5-Instruct 7B.
-“All/Err” denotes the total number of reasoning steps and the step which a fault was injected.
-“Type” indicates the category of the injected error.
-“Ori Acc” is the model’s accuracy before error-injection.
-“Fin Acc” is the accuracy after injection.
-“Accu Drop(abs.
-points)” quantifies the relative decrease in accuracy.
-All/Err Type Ori Acc Fin Acc Accu Drop
+9–11 17 20 85.00
+5.5 Ablation Study
+To rigorously evaluate the contribution of each component within ASCoT, we conducted a multi-level ablation study on the GSM8K dataset.
+We first validated the necessity of the core modules: IRM and AVM.
+Subsequently, we performed a fine-grained analysis of the MSCE module to disentangle the efficacy of intrinsic versus extrinsic correction strategies.
+Our results detailed in Appendix F, confirm that the synergistic integration of position-aware verification and dual-path correction is essential for achieving state-of-the-art performance while minimizing token usage.
+Table 4: Error-injection results on GSM8K using Qwen2.5-Instruct 7B. “All/Err” denotes the total steps vs. injected step. “Drop” denotes the accuracy drop.
+All/Err Type Ori Acc Fin Acc Drop
 4/2 numeric 95.75 81.11 14.64
@@
 3/3 96.19 39.41 56.78
-5.2. CoT Compression and Pruning
-The verbosity of CoT Wingate et al.
-(2022), while beneficial for reasoning, leads to high inference latency and computational costs.
-This has spurred research into CoT compression and pruning Chuang et al.
-(2024); Xuefei et al.
-(2023), which aims to eliminate redundant information without degrading reasoning accuracy.
-This issue is often termed the “overthinking phenomenon,” Kumar et al.
-(2025); Chen et al.
-(2024a) where models generate unnecessarily long chains that can even introduce errors.
-To address this bottleneck, researchers have explored various techniques.
-Token-level pruning Fu et al.
-(2024); Ma et al.
-(2023), uses a semantic importance metric to identify and remove less critical tokens from the reasoning chain, training the model to learn shortcuts between essential concepts.
-Other approaches operate at a more structural level.
-Prune-on-Logic Liu et al.
-(2025) transforms a CoT into a logic graph and prunes entire reasoning steps based on a loss importance score.
-Simpler methods, such as direct truncation of the reasoning chain Wang et al.
-(2024), have also been shown to be surprisingly effective, suggesting that complete reasoning paths are not always necessary.
-5.3. Self-correction Mechanism
-The capacity for self-correction is a critical frontier Pan et al.
-(2023) for enhancing the reliability of LLMs.
-The general paradigm Kamoi et al.
-(2024); Yang et al.
-(2025) involves a multi-stage process: generating an initial response, producing feedback on that response, and then refining it.
-Approaches diverge based on the source of the feedback.
-Intrinsic self-correction Li et al.
-(2024); Liu et al.
-(2024c) relies on the model’s own capabilities to critique and improve its output without external tools Cao et al.
-(2025); Tan et al.
-(2025).
-However, a key challenge is that models often fail to identify their own errors, a phenomenon attributed to a lack of error-correction examples in training data Liu et al.
-(2024b); Anantaprayoon et al.
-(2025).
-Extrinsic self-correction incorporates feedback from external sources.
-This involves using a verifier model to assess the correctness of each step Kim et al.
-(2025), query external knowledge, or check the consistency across outputs generated from multiple perspectives.
-6. Limitations
-While ASCoT demonstrates promising performance, we acknowledge that there still exist several limitations that define avenues for future research.
-The efficacy of the framework is inherently dependent on the self-evaluation capabilities of its base Large Language Model (LLM); if a model cannot fundamentally recognize its own fallacies, the utility of the AVM and MSCE modules is consequently constrained.
-Furthermore, our current evaluation is focused on numeric and symbolic errors within the specific domain of mathematical reasoning.
-Therefore, the generalization of ASCoT to other error modalities—such as factual, commonsense, or complex logical errors—and its transferability to diverse long-chain reasoning tasks like code generation, legal analysis, or scientific summarization, remain important open questions requiring further study.
-7. Conclusion
-In this work, we challenge the conventional “cascading failure” hypothesis in CoT reasoning.
-Through systematic error-injection experiments, we identify and quantify “Late-Stage Fragility”: a counter-intuitive phenomenon where errors introduced near the end of a reasoning chain are significantly more detrimental than those at the beginning.
-To counteract this specific vulnerability, we introduce the Adaptive self-correction Chain-of-Thought (ASCoT) method, which is a targeted system that uses an adaptive verification manager to precisely identify and a multi-perspective engine to robustly correct these high-impact, late-stage errors.
-Our extensive experiments on the GSM8K and MATH benchmarks demonstrate that ASCoT achieves outstanding accuracy, outperforming strong baselines including the computationally expensive self-consistency.
-These results highlight the effectiveness of ASCoT in addressing late-stage fragility, validating the importance of diagnosing specific failure modes to engineer more effective solutions.
-This research advocates for a shift from uniform verification to adaptive, position-aware strategies.
-We hope our investigations pave the way for future advancements in correction mechanisms and enhance the reliability, robustness, and efficiency of LLM reasoning.
+6 Conclusion
+In this work, we critically re-examine the dynamics of error propagation in CoT reasoning, identifying a counter-intuitive phenomenon termed Late-Stage Fragility.
+Our systematic analysis reveals that contrary to the prevailing cascading failure assumption, errors introduced in the terminal stages of reasoning impose a disproportionately higher penalty on the final outcome.
+To address this vulnerability, we introduce ASCoT (Adaptive Self-Correction Chain-of-Thought).
+By integrating an Intelligent Routing Mechanism with a position-aware Adaptive Verification Manager, ASCoT strategically reallocates computational resources to verify high risk, late-stage steps.
+Extensive experiments on GSM8K and MATH benchmarks demonstrate that ASCoT achieves a superior trade-off between reasoning fidelity and computational efficiency, outperforming strong baselines including Self-Consistency.
+Ultimately, our findings advocate for a paradigm shift from uniform verification to adaptive, context-sensitive strategies, establishing a new foundation for reliable and resource-efficient LLM reasoning.
+Limitations
+While ASCoT effectively mitigates Late-Stage Fragility, its performance naturally correlates with the base model’s verification capabilities.
+Additionally, our current study primarily targets mathematical reasoning focusing on numeric and symbolic errors.
+Extending the framework to broader domains ($e.g.,$ code generation) and diverse modalities remains a promising avenue for future exploration.
 Appendix A Appendix
@@
 Additionally, the document discusses the model’s Late-Stage Fragility and outlines current limitations and future directions.
-A.1. Experimental Setup
+Appendix B Algorithmic Details
+Algorithm 1 formally outlines the inference pipeline of the proposed ASCoT framework.
+The process initiates with the Intelligent Routing Mechanism (IRM), which semantically prunes the initial chain $\mathcal{C}_{\text{init}}$ to filter redundant tokens, yielding a compressed sequence $\mathcal{C}_{\text{pruned}}$.
+Subsequently, the system iterates through each reasoning step $t_{k}$.
+At each juncture, the Adaptive Verification Manager (AVM) assesses the step against the accumulated context $S_{\text{ctx}}$, computing a risk score based on both content quality and positional impact.
+If this score exceeds the predefined threshold $\tau$, the step is flagged as fragile.
+In such cases, the Multi-Perspective Self-Correction Engine (MSCE) is triggered to generate a rectified step $t_{k}^{\text{corr}}$ via dual-path correction.
+Conversely, low-risk steps are retained directly.
+This selective intervention ensures that computational resources are reallocated specifically to fix late-stage vulnerabilities while maintaining overall efficiency.
+Algorithm 1 ASCoT: Adaptive Self-Correction Chain-of-Thought
+1:Initial CoT $\mathcal{C}_{\text{init}}$; Compression ratio $\gamma$; Risk threshold $\tau$
+2:Corrected CoT $\mathcal{C}_{\text{final}}$
+3:// Stage 1: Semantic Pruning via IRM
+[…]
+5:// Stage 2 & 3: Adaptive Verification and Correction
+[…]
+7:for each step $t_{k}$ in $\mathcal{C}_{\text{pruned}}$ do
+8:  // Calculate risk via AVM (Logic, Fact, Clarity, Position)
+9:  $R_{k}\leftarrow\mathrm{ComputeRiskScore}(t_{k},S_{\text{ctx}})$
+10:  if $R_{k}>\tau$ then
+11:   // High risk detected: Trigger MSCE correction
+12:   $t_{k}^{\text{corr}}\leftarrow\mathrm{DualPathCorrection}(t_{k},S_{\text{ctx}})$
+13:   $\mathrm{Append}(t_{k}^{\text{corr}},\mathcal{C}_{\text{final}})$ $\triangleright$ Use corrected step
+14:   $S_{\text{ctx}}\leftarrow\mathrm{UpdateContext}(S_{\text{ctx}},t_{k}^{\text{corr}})$
+15:  else
+16:   $\mathrm{Append}(t_{k},\mathcal{C}_{\text{final}})$ $\triangleright$ Keep original step
+17:   $S_{\text{ctx}}\leftarrow\mathrm{UpdateContext}(S_{\text{ctx}},t_{k})$
+18:  end if
+19:end for
+20:// Final Output Generation
+21:return $\mathcal{C}_{\text{final}}$
+Appendix C Qualitative Analysis of Late-Stage Fragility
+To visually demonstrate the Late-Stage Fragility phenomenon discussed in Section 4.3, we provide a comparative case study in Figure 6.
+The example contrasts the model’s robust recovery from early-stage perturbations against its failure to correct identical errors in the final reasoning steps.
+Figure 6: An illustration of Late-Stage Fragility.
+The model recovers from an early-stage calculation error, but fails when an identical type of error is introduced in the final step.
+The tokens that occur related errors are highlighted in orange.
+Appendix D Experimental Setup
 To clearly elucidate the algorithmic pipeline of the ASCoT methodology, we provide its pseudocode.
@@
 For each step $t_{k}$, the AVM computes a Risk Score $R(t_{k})$, which integrates a multi-dimensional quality assessment $Q(t_{k})$ with a Positional Impact score $I(k)$.
-If this score exceeds a predefined threshold $\tau$, the step is flagged as high-risk and is routed to the MSCE module for remediation.
+If this score exceeds a predefined threshold $\tau$, the step is flagged as high risk and is routed to the MSCE module for remediation.
 The MSCE employs a dual-path strategy, encompassing intrinsic reflection and extrinsic independent generation to correct the identified error, yielding the final, refined $CoT_{final}$.
@@
 During inference, the maximum number of tokens set to 512 for GSM8K and 1024 for MATH.
-A.2. Error‑Injection Experiments
-A.2.1. Error Simulation
+Appendix E Error‑Injection Experiments
+E.1 Error Simulation
 To probe the model’s resilience in a realistic manner, our error-injection methodology does not introduce arbitrary flaws.
@@
 Following this injection, the model was prompted to resume the reasoning process from that point, and the accuracy of the final answer was recorded to precisely measure the positional impact of the error.
-A.2.2. Distribution of Results
-Table 5 illustrates the proportion of total steps across all problems.
-We can conclude from the chart that a significant majority of problems are concentrated in the 2-5 step range, with 3-step and 4-step problems being the most prevalent.
-When injecting a numerical error, the accuracy drops are relatively uniform across positions: at step 2 yields a drop of 59.92%, at step 3 a drop of 60.47%, and at step 4 a drop of 61.70%a miscalculation.
-In contrast, symbolic errors induce more severe degradation, and this severity escalates with later injection points.
-Replacement of an operator in the first error step (2/2) reduces the accuracy by 69.67%, while an error in the final step (4/4) inflicts a reduction of 80.80%, with symbolic faults becoming increasingly catastrophic when introduced deeper into the chain.
 Infer Steps Counts Accuracy
@@
 9 7 57.14
-Table 5. Performance analysis by reasoning step count on the GSM8K dataset, illustrating both the frequency of problems per step count and the corresponding model accuracy.
-A.3. Analysis of Late‑Stage Fragility
-Early-stage errors often activate the model’s latent self-correction capability, wherein subsequent reasoning steps can reinterpret and compensate for previous mistakes, leading to a correct final answer.
+Table 5: Performance analysis by reasoning step count on the GSM8K dataset, illustrating both the frequency of problems per step count and the corresponding model accuracy.
+E.2 Distribution of Results
+Table  5 illustrates the proportion of total steps across all problems.
+We can conclude from the chart that a significant majority of problems are concentrated in the 2-5 step range, with 3-step and 4-step problems being the most prevalent.
+When injecting a numerical error, the accuracy drops are relatively uniform across positions: at step 2 yields a drop of 59.92%, at step 3 a drop of 60.47%, and at step 4 a drop of 61.70%a miscalculation.
+In contrast, symbolic errors induce more severe degradation, and this severity escalates with later injection points.
+Replacement of an operator in the first error step (2/2) reduces the accuracy by 69.67%, while an error in the final step (4/4) inflicts a reduction of 80.80%, with symbolic faults becoming increasingly catastrophic when introduced deeper into the chain.
+Analysis of Late‑Stage Fragility Early-stage errors often activate the model’s latent self-correction capability, wherein subsequent reasoning steps can reinterpret and compensate for previous mistakes, leading to a correct final answer.
 This is facilitated by the model maintaining higher semantic entropy and contextual flexibility during the initial stages of reasoning.
@@
 We refer to this phenomenon as semantic commitment, a state in which the model’s outputs become rigidly anchored to previously generated content.
-At this point, even minor numerical or symbolic errors near the conclusion are less likely to be detected or corrected, resulting in disproportionately high failure rates.
+At this point, even minor numerical or symbolic errors near the conclusion are less likely to be detected or corrected, resulting in high failure rates.
 Our findings do not aim to refute the cascading failure hypothesis entirely, but rather to refine and contextualize it.
@@
 This indicates that cascading failures do occur, but that Late-Stage Fragility constitutes an additional failure mode not previously emphasized.
-Figure 7. Example generations by ASCoT, using Qwen2.5-Instruct 7B.
-A.4. Generation Examples
+Figure 7: Example generations by ASCoT, using Qwen2.5-Instruct 7B.
+Appendix F Detailed Ablation Studies
+In this section, we present a series of ablation experiments to validate the design choices of ASCoT. All experiments were conducted using the LLaMA-3.1-8B model on the GSM8K benchmark.
+F.1 Impact of Core Components
+We first investigate the necessity of the two fundamental stages: semantic pruning (IRM) and the correction pipeline (AVM+MSCE). Table 6 summarizes the results.
+Table 6: Ablation of core components. “w/o IRM” denotes removing the pruning stage; w/o Correction denotes using only pruning without any verification or correction.
+Configuration Acc. (%) Tokens ActRatio
+ASCoT (Full) 86.9 214.24 1.00
+w/o IRM 87.1 325.40 1.52
+w/o Correction 82.5 198.12 0.92
+As shown in Table 6, removing the IRM leads to a massive increase in token usage (+52%) with negligible accuracy gain (+0.2%), validating IRM’s critical role in efficiency.
+Conversely, removing the correction pipeline (w/o Correction) causes accuracy to drop significantly to 82.5%, underscoring that semantic pruning alone induces fragility that must be compensated for by our robust verification mechanism.
+F.2 Efficacy of Position-Aware Verification
+A key innovation of ASCoT is the integration of the Positional Impact Score within the AVM module to address Late-Stage Fragility. Table 7 compares our strategy against a uniform risk assessment.
+Table 7: Ablation of the AVM strategy. We compare our position-aware scoring against a uniform weighting scheme.
+AVM Strategy Acc. (%) $\Delta$ Acc.
+Position-Aware (Ours) 86.9
+Uniform Weights 84.3 -2.6
+The results in Table 7 reveal a 2.6% performance degradation when position awareness is removed.
+This empirical evidence supports our hypothesis that errors in later reasoning steps are more detrimental.
+By assigning higher risk weights to these terminal steps, ASCoT effectively reallocates verification resources to where they are most needed.
+F.3 Analysis of Dual-Path Correction
+Finally, we dissect the Multi-Perspective Self-Correction Engine (MSCE) to evaluate the contribution of its Dual-Path (Intrinsic + Extrinsic) strategy compared to single-source baselines.
+Table 8: Ablation of the MSCE correction strategy. “Intrinsic” relies on self-reflection; “Extrinsic” uses independent regeneration.
+Correction Strategy Acc. (%) $\Delta$ Acc.
+Dual-Path (Ours) 86.9
+Extrinsic Only 86.1 -0.8
+Intrinsic Only 85.0 -1.9
+Table 8 demonstrates that the Dual-Path strategy outperforms individual approaches.
+Relying solely on Intrinsic correction yields the lowest accuracy (85.0%), likely because the model struggles to escape its own erroneous reasoning context.
+Extrinsic correction performs better (86.1%) but is surpassed by the synergistic combination of both, which leverages the strengths of context-aware reflection and objective regeneration.
+Appendix G Generation Examples
 To provide a qualitative illustration of our method’s end-to-end functionality, we present representative examples of reasoning chains after processing by the ASCoT pipeline.
-The Figure 7 displays the corrected output after being processed by ASCoT.
+Figure 7 displays the corrected output after being processed by ASCoT.
 This transformation from a flawed to a correct reasoning chain highlights the core capabilities of our system.
-It demonstrates the AVM module successfully identifying a high-risk step by analyzing its content quality and positional impact, followed by the MSCE module executing a targeted repair.
-This example visually confirm ASCoT’s ability to structure reasoning but, more importantly, to dynamically detect and rectify internal inconsistencies, ultimately leading to a more trustworthy final answer.
-A.5. Future Work
+It demonstrates the AVM module successfully identifying a high risk step by analyzing its content quality and positional impact, followed by the MSCE module executing a targeted repair.
+This example confirms ASCoT’s ability to structure reasoning but, more importantly, to dynamically detect and rectify internal inconsistencies, ultimately leading to a more trustworthy final answer.
+Appendix H Future Work
 We propose several promising future directions.